### v2

# ESC: Efficient Speech Coding with Cross-Scale Residual Vector Quantized Transformers

Abstract
Neural speech codecs aim to compress input signals into minimal bits while maintaining content quality in a low-latency manner. However, existing codecs often trade computational complexity for reconstruction performance. These codecs primarily use convolutional blocks for feature transformation layers, which are not inherently suited for capturing the local redundancies in speech signals. To compensate, they require either adversarial discriminators or a large number of model parameters to enhance audio quality. In response to these challenges, we introduce the Efficient Speech Codec (ESC), a lightweight, parameter-efficient speech codec based on a cross-scale residual vector quantization scheme and transformers. Our model employs mirrored hierarchical window transformer blocks and performs step-wise decoding from coarse-to-fine feature representations. To enhance bitrate efficiency, we propose a novel combination of vector quantization techniques along with a pre-training paradigm. Extensive experiments demonstrate that ESC can achieve high-fidelity speech reconstruction with significantly lower complexity, making it a promising alternative to existing convolutional audio codecs [^1] Our code and pretrained models are available here..
ESC: Efficient Speech Coding with Cross-Scale Residual Vector Quantized Transformers
Yuzhe Gu1,2, Enmao Diao2 1University of Pennsylvania, Philadelphia, PA 2Duke University, Durham, NC tracygu@seas.upenn.eduenmao.diao@duke.edu
1 Introduction
Recent advancements in deep learning have demonstrated the superiority of neural speech codecs over traditional ones, which rely on complex expert design and psycho-acoustic knowledge Valin et al. (2012); Dietz et al. (2015). Early efforts integrating deep generative models such as WaveNet Oord et al. (2016) and SampleRNN Mehri et al. (2017) into audio codecs have yielded promising results. These models serve as powerful decoders, synthesizing high-quality speech audio based on the intermediate representations generated by traditional codecs Kleijn et al. (2018); Klejsa et al. (2019). However, their auto-regressive decoding nature often introduces significant inference latency, limiting their practical application.
Alternatively, some end-to-end neural audio codecs build upon the vector quantization (VQ) network first introduced by Van Den Oord et al. (2017). VQ networks, which consist of trainable codebooks, can discretize continuous input vectors by assigning them to the nearest neighboring code-vectors, making them suitable for both generation and compression tasks. Based on this framework, existing VQ codecs Zeghidour et al. (2021); Défossez et al. (2023); Kumar et al. (2023) typically follow a three-stage architecture with a convolutional encoder and decoder interleaved with a residual vector quantization (RVQ) module Vasuki and Vanathi (2006) at the bottleneck. The encoder and decoder perform downsampling and upsampling of audio signals in the time-domain to produce hierarchical representations. RVQ is an enhanced quantization scheme where a series of VQs recursively quantize the residuals from previous stages. Additionally, these codecs utilize generative adversarial networks (GANs) to remove artifacts and produce high-fidelity audio reconstructions. Significant effort has been dedicated to developing effective audio discriminators that leverage multi-scale features of waveforms and spectrograms to improve the generated audio during training Esser et al. (2021); Kumar et al. (2019); Kong et al. (2020). Currently, VQ-based codecs have demonstrated remarkable capabilities in audio reconstruction even at ultra-low bitrates.
Despite these advantages, we find that convolutional VQ-based codecs heavily rely on powerful GAN discriminators to produce high-quality audio, introducing additional optimization challenges due to adversarial training. Moreover, convolutional codecs tend to face computational limitations, as they require a large number of parameters to achieve a good trade-off between compression rate and reconstruction quality. To address these issues, our work aims to develop a more parameter-efficient speech codec by implementing the following architectural improvements: 1) Replacing convolutional layers with efficient Swin-Transformer Blocks (STBs) Liu et al. (2021), which have greater capacity for modeling acoustic features; 2) Utilizing the cross-scale residual vector quantization (CS-RVQ) scheme Jiang et al. (2022a) instead of RVQ, extending quantization from a fixed level to multiple levels.
In addition, training VQ-based codecs encounters a significant obstacle: a fraction of the vector quantizer’s codebook is often underutilized for representing input vectors, leading to a phenomenon known as codebook collapse. This issue is frequently observed in image generation tasks Takida et al. (2022); Zhang et al. (2023); Huh et al. (2023). To address this in the context of speech compression, we propose combining product vector quantization (PVQ) Baevski et al. (2019), latent dimension reduction Yu et al. (2022), and Euclidean normalization Łańcucki et al. (2020) to enhance codebook utilization. Additionally, we introduce a learning paradigm to facilate codebook optimization. It incorporates a pre-training stage during which VQs are freezed and trained afterwards.
In summary, the contributions of our work are as follows:
  • We introduce ESC, a lightweight efficient speech codec that integrates cross-scale residual vector quantization (CS-RVQ) and mirrored hierarchical Swin-Transformer blocks, achieving a superior trade-off between reconstruction performance and codec complexity compared to current state-of-the-art models.
  • We propose a novel combination of vector quantization techniques along with a pre-training paradigm that includes a warm start for the encoder and decoder, aimed at mitigating codebook collapse and enhancing bitrate efficiency.
  • Our extensive experiments demonstrate that transformers and CS-RVQ, the core components of ESC, are more efficient backbones for speech compression than convolutions and RVQ.
2 Related Work
2.1 VQ-based Audio Codecs
Notable examples in recent VQ-based audio codecs include SoundStream Zeghidour et al. (2021), EnCodec Défossez et al. (2023), and Descript’s Audio Codec (DAC) Kumar et al. (2023). SoundStream is recognized as the first universal neural codec designed for various audio types. EnCodec enhances compression rates by integrating a lightweight transformer language model at the discrete latent space and explores a streaming architecture. Following similar methodologies, Kumar et al. (2023) identify the side effects of quantization dropout, a technique for bitrate scalability, and demonstrate the superiority of a periodic inductive bias function over other common activation functions in audio signal modeling gil Lee et al. (2023); Ziyin et al. (2020). These codecs process audio signals directly in the time domain and are thus categorized as time-domain codecs. Meanwhile, frequency-domain codecs, such as TFNet Jiang et al. (2022b) and our proposed ESC, leverage the invertible short-time Fourier transform (STFT) to convert time-domain audio into spectrograms. The encoder and decoder thus process these two-dimensional spectral features directly. Among recent neural codecs, DAC achieves state-of-the-art compression ratios and audio reconstruction quality, although its computation bottleneck is ignored.
2.2 Swin Transformers
Vision Transformers (ViTs) Dosovitskiy et al. (2020) have outperformed convolutional neural networks (CNNs) in various image processing tasks due to their superior ability to capture complex patterns. The Swin Transformer Liu et al. (2021), a notable variant, enhances this capability through a hierarchical approach that leverages window attention mechanisms. It can scale efficiently to high-resolution signals while maintaining computational efficiency. In image compression, Swin Transformers have demonstrated exceptional performance. Studies by Zhu et al. (2021) and Zou et al. (2022) have shown that they outperform CNNs in modeling spatial hierarchies and long-range dependencies. The attention mechanisms enable the accurate preservation of essential details and textures, even at lower bitrates. These capabilities highlight the potential of transformers for applications beyond image compression, suggesting their suitability for the field of speech compression as well.
2.3 Vector Quantization
Proposed in Vector Quantized Variational Autoencoders (VQ-VAEs) (Van Den Oord et al., 2017), vector quantization (VQ) serves as a trainable layer that deterministically quantizes the encoded latent variables to their nearest neighbors in an embedding codebook. A vector quantization layer $Q(\cdot\ ;\mathcal{C})$ is parameterized by a codebook collection of continuous vectors $\mathcal{C}={\bm{c}_{1},...,\bm{c}_{K}}$. Each vector is referred to as a codeword, and its index is known as a code. The layer quantizes a vector $\bm{z}_{e}\in\mathbb{R}^{d}$ to be $\bm{z}_{q}\in\mathbb{R}^{d}$ by selecting the Euclidean nearest codeword $\bm{c}_{k}$ from the codebook $\mathcal{C}$, i.e.,
$\bm{z}_{q}:=\bm{c}_{k}=\operatorname*{arg\,min}_{\bm{c}_{j}\in\mathcal{C}}||\bm{z}_{e}-\bm{c}_{j}||_{2}^{2}.$ (1)
For notation convenience, we denote the vector quantization function as $\bm{z}_{e}\xrightarrow[]{Q(\cdot;\mathcal{C})}(\bm{z}_{q},\tilde{z}_{q})$, where $\tilde{z}_{q}:=k$ is the corresponding discrete code for the nearest codeword. During compression, the encoding process outputs the discrete index $\tilde{z}_{q}$, which is stored with a $\log_{2}K$ bit budget. The decoding process starts by retrieving the continuous vector $\bm{z}_{q}$ from the codebook using the encoded index $\tilde{z}_{q}$. The VQ function $Q(\cdot;\mathcal{C})$ is non-differentiable due to the $\operatorname*{arg\,min}$ operator. Common strategies use a straight-through estimator (STE) Bengio et al. (2013) to bypass this in back-propagation. In other words, the gradient component $\frac{\partial\bm{z}_{q}}{\partial\bm{z}e}$ is eliminated, and as compensation, a codebook loss and a commitment loss are proposed to pull together the codewords and the encoded latents:
$\mathcal{L}_{vq}=||\text{sg}(\bm{z}_{e})-\bm{z}_{q}||_{2}^{2}+\beta||\bm{z}_{e}-\text{sg}(\bm{z}_{q})||_{2}^{2}.$ (2)
Here $\text{sg}(\cdot)$ denotes the stop gradient operator. The first term updates the codebook with an $l_{2}$ error, moving the codewords towards the encoder outputs. The second term ensures $\bm{z}_{e}$ can commit to the embedding without growing arbitrarily. The scalar $\beta$ is used to balance the importance of updating the codebook and encoder.
2.4 Codebook Collapse
Straight-through estimators (STEs) often lead to significant issues, most notably the issue of codebook collapse Vuong et al. (2023). In a recent study, Huh et al. (2023) offer a straightforward explanation, attributing it to an internal codebook covariate shift during the training phase. Frequent adjustments in encoder representations cause misalignment with VQ codebooks, resulting in only a subset of codewords being updated. Consequently, VQ layers are prone to divergence, often ending up with a significant number of dead vectors. Various strategies have been proposed in generative modeling tasks to address this problem, including the use of stochastic VQs Takida et al. (2022); Zhang et al. (2023), self-annealing soft-to-hard quantization approaches Agustsson et al. (2017), re-initializing codewords with K-means centroids every few iterations Łańcucki et al. (2020); Dhariwal et al. (2020), and avoiding the issues of STEs with finite scalar quantizers Mentzer et al. (2024). In audio compression, Kumar et al. (2023) attempt to mitigate codebook collapse by reducing the dimensions of codewords Yu et al. (2022) and normalizing them within a Euclidean ball Łańcucki et al. (2020).
Figure 1: The framework of ESC: Input speech is transformed to a complex STFT $\mathcal{X}$ and linearly embedded into patches. Encoder STBs iteratively halve the frequency resolution and produce hierarchical feature representations. Mirrored decoder STBs iteratively recover the frequency resolution by leveraging coarse-to-fine encoded features. The intermediate residuals between encoder and decoder hidden states are progressively quantized to refine decoding. The entire network is solely composed of efficient transformer blocks. The figure displays a scenario when the deepest $3$ of $n+1$ total bitstreams (solid lines) are transmitted, with others left inactive.
3 Efficient Speech Codec (ESC)
3.1 Overall Architecture
As illustrated in Figure 1, our codec operates on the complex spectrum $\mathcal{X}\in\mathbb{R}^{2\times F\times T}$ derived from the Short-Time Fourier Transform (STFT) of a speech signal. Here, the real and imaginary components of $\mathcal{X}$ are treated as separate channels. The model consists of a series of mirrored transformer encoder and decoder layers, each performing downsampling or upsampling to generate coarse and fine representations, as detailed in Section 3.3. Starting from the quantized latents at the bottleneck VQ, the decoder progressively reconstructs the original spectrum by leveraging multi-level residuals between the intermediate features of the encoder and decoder. This cross-scale decoding mechanism is explained in further detail in Section 3.4. Finally, the reconstructed spectrum $\hat{\mathcal{X}}$ is transformed back into the time domain through the inverse STFT.
3.2 Notations
We first define some architectural notations for clarity. We represent the encoder and decoder functions $F_{\phi}(\cdot)$ and $G_{\psi}(\cdot)$, each as a composition of individual layer functions $f_{\phi_{1}},\ldots,f_{\phi_{n}}$ and $g_{\psi_{1}},\ldots,g_{\psi_{n}}$, respectively. We use $\mathcal{Z}$ to represent spectrum feature and $\bm{z}$ to denote a feature frame. Specifically, we denote $\mathcal{Z}_{e_{i}}$ as the hidden feature after $i$ encoder layers, and $\mathcal{Z}_{q_{i}}$ as the corresponding $i$-th decoder feature. Formally, this is expressed as:
$\mathcal{Z}_{e_{i}}=f_{\phi_{i}}\circ...\circ f_{\phi_{1}}\circ\mathcal{Z}_{e_{0}}$ (3)
$\mathcal{Z}_{q_{i}}=g_{\psi_{i}}\circ...\circ g_{\psi_{1}}\circ\mathcal{Z}_{q_{0}},$ (4)
where $\mathcal{Z}_{e_{0}}$ is the original input feature and $\mathcal{Z}_{q_{0}}$ is the latent representation at the bottleneck.
3.3 Transformer Encoder and Decoder
To effectively capture redundancies within audio signals, we replace convolutional layers with hierarchical Swin Transformer blocks (STBs) and their extended decoding counterparts.
Patchify. The encoder begins with a linear patchify module, where the complex spectrum $\mathcal{X}$ is divided into small patches and linearly projected:
[…]
Here, the patch size across the frequency and temporal dimensions is $(\frac{F}{H_{0}},\frac{T}{W_{0}})$. This step reduces the input resolution to alleviate the computational burden on transformers. At the end of the decoder, a symmetric de-patchify module reshapes the decoded patch embedding $\mathcal{Z}_{q_{n}}$ and linearly projects it back to produce a recovered spectrum $\hat{\mathcal{X}}$.
Swin Transformer blocks. STBs in both the encoder and decoder employ window-based multi-head self-attention (W-MSA), partitioning spectrum features into smaller windows and computing attention in parallel within each window. This approach enables more efficient computation compared to vanilla attention mechanisms. To ensure connections between windows, STBs cascade two interleaved W-MSAs, with the outputs of the first being shifted for the second. This design allows STBs to capture local and global feature dependencies both effectively and efficiently.
Downsampling and upsampling. Our codec maintains temporal resolution while scaling frequency resolution to equalize bitrates across different bitstreams. To achieve this, we modify the original patch merging/splitting modules with a single-dimensional pixel unshuffle/shuffle module Shi et al. (2016) along the frequency dimension. During encoder downsampling, an intermediate encoder spectrum feature $\mathcal{Z}_{e_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is first reshaped and then projected as follows:
[…]
where $v$ is the scaling factor. The upsampling process mirrors this operation in reverse. An intermediate decoder feature $\mathcal{Z}_{q_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is up-projected and then reshaped, resulting in an up-scaled frequency resolution:
[…]
Thus, the transformer encoder and decoder layers are mirrored, creating symmetric and hierarchical representations of the input audio spectrum. With these backbones, ESC is a fully transformer-based codec without any convolutional modules.
3.4 Cross-Scale Residual Vector Quantization
For efficient modeling of audio signals, ESC employs multi-scale features through a coarse-to-fine decoding process to generate bitstreams at different levels. Our codec integrates the more intuitive cross-scale residual vector quantization (CS-RVQ) structure proposed by Jiang et al. (2022a), eliminating the need for additional networks to fuse encoder and decoder features. As depicted in Algorithm 1, Algorithm 2 and Figure 1, the decoding process is conditioned on encoder features from various resolution scales. This distinguishes it from the commonly used residual vector quantization scheme that operates only at the bottom scale, which often overlooks low-level information.
Encoding. The encoding process begins with the encoder function $F_{\phi}(\cdot)$, creating multi-scale encoder features $\bm{z}_{e_{1}},...,\bm{z}_{e_{n}}$. $\bm{z}_{e_{n}}$ is first quantized by the bottleneck quantizer $Q_{0}$ to form the lowest bitstream. This represents the simplest case when $s$ is set to 1, and CS-RVQ reduces to a fixed-scale VQ at the bottleneck. At higher bitstreams, the residual between symmetric encoder and decoder features with higher resolutions, $\bm{z}_{e_{n-i+1}}-\bm{z}_{q_{i-1}}$, is quantized by $Q_{i}$. This quantized residual $\bm{q}_{i}$ is added back to $\bm{z}_{q_{i-1}}$ and decoded by the subsequent decoder layer $g_{\phi_{i}}(\cdot)$, producing the next decoder feature $\bm{z}_{q_{i}}$. Recursively, residuals at higher resolutions are progressively quantized, forming the remaining bitstreams. This mechanism enables multi-scale learning and allows the decoder layers to incrementally reduce quantization errors by conditioning on encoder features. When $s>2$, the encoding process of CS-RVQ requires $s-2$ additional decoder layers to produce residuals at higher levels. After encoding, $\bm{z}_{e_{0}}$ is compressed to be multi-level codes $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},\ldots,\tilde{z}_{q_{s-1}}$.
Decoding. The decoding process starts by retrieving the quantized latent from the bottom VQ using the code index $\tilde{z}_{q_{0}}$, which provides the original decoder input $\bm{z}_{q_{0}}$. At higher levels, $\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$ are iteratively used to retrieve codewords, resulting in multi-scale quantized residuals $\bm{q}_{1},...,\bm{q}_{s-1}$. Each quantized residual $\bm{q}_{i}$ is added back to the corresponding decoder feature $\bm{z}_{q_{i-1}}$ to improve the decoding process. Starting from the $s$-th decoder layer, there are no quantized residuals, and the layers perform vanilla decoding. Finally, the recovered frame spectrum $\bm{z}_{q_{n}}$ is obtained, which conditions on $s-1$ residual features.
Training. During the training of CS-RVQ codecs, the encoding and decoding processes are combined to form a complete forward pass. To enable bitrate scalability, we sample the number of bitstreams $s\sim\text{Uniform}\{1,\ldots,n\}$ at a rate $p$ within each training batch. $p$ is a hyperparameter that balances the reconstruction quality at different bitrates, as proposed by Kumar et al. (2023).
Algorithm 1 CS-RVQ Encoding
1:Patch embedding frame $\bm{z}_{e_{0}}\in\mathbb{R}^{C_{0}\times H_{0}}$, encoder $F_{\phi}(\cdot)$, decoder $G_{\psi}(\cdot)$, vector quantizers $Q_{0},Q_{1},...,Q_{s-1}$, number of bitstreams $s$
2:$\bm{z}_{e_{1}},...,\bm{z}_{e_{n}}\leftarrow F_{\phi}(\bm{z}_{e_{0}})$triangleright Encoder forward pass
3:$\bm{z}_{q_{0}},\tilde{z}_{q_{0}}\leftarrow Q_{0}(\bm{z}_{e_{n}})$triangleright Bottleneck VQ
4:for $i=1\dots s-2$ do
5:     $\bm{q}_{i},\tilde{z}_{{q}_{i}}\leftarrow Q_{i}(\bm{z}_{e_{n-i+1}}-\bm{z}_{{q}_{i-1}})$
6:     $\bm{z}_{{q}_{i}}\leftarrow g_{\psi_{i}}(\bm{z}_{{q}_{i-1}}+\bm{q}_{i})$
7:end $\triangleright$ Encoding requires $s-2$ decoder layers
8:if $s>1$ then
9:     $\bm{q}_{s-1},\tilde{z}_{{q}_{s-1}}\leftarrow Q_{i}(\bm{z}_{e_{n-s+2}}\mathrel{-}\bm{z}_{{q}_{s-2}})$
10:end if
11:return $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$
Algorithm 2 CS-RVQ Decoding
1:Codes $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$, decoder $G_{\psi}(\cdot)$, vector quantizers $Q_{0},Q_{1},...,Q_{s-1}$
2:$\bm{z}_{q_{0}}\xleftarrow[]{Q_{0}}\tilde{z}_{q_{0}}$triangleright Retrieve codewords from VQ
3:for $i=1\dots s-1$ do
4:     $\bm{q}_{i}\xleftarrow[]{Q_{i}}\tilde{z}_{q_{i}}$
5:     $\bm{z}_{{q}_{i}}\leftarrow g_{\psi_{i}}(\bm{z}_{{q}_{i-1}}+\bm{q}_{i})$
6:end $\triangleright$ Decoding conditions on residuals
7:for $i=s\dots n$ do
8:     $\bm{z}_{{q}_{i}}\leftarrow g_{\psi_{i}}(\bm{z}_{{q}_{i-1}})$
9:end $\triangleright$ Continue with vanilla decoding
10:return $\bm{z}_{q_{n}}$
3.5 Mitigating Codebook Collapse
During quantization, the spectrum frame feature needs to be flattened, merging the frequency and channel dimensions. This results in large hidden dimensions for VQ, increasing the possibility of codebook collapse during training.
Vector quantization setups. To effectively optimize the codebooks, we modify vanilla VQs using a combination of product vector quantization and vector dimension reduction at each bitstream. Specifically, a frame feature $\bm{z}_{e_{i}}\in\mathbb{R}^{C_{i}\times H_{i}}$ is flattened into a $d$-dimensional vector before quantization. It is then split into a collection of $l$ sub-vectors. Each sub-vector $\bm{z}_{e_{i}}^{(m)}$ is down-projected and quantized by an individual codebook $\mathcal{C}_{m}$ to be ${\bm{z}}_{q_{i}}^{(m)}$:
$\bm{z}_{e_{i}}\equiv\{\bm{z}_{e_{i}}^{(m)}\mid\bm{z}_{e_{i}}^{(m)}\in\mathbb{R}^{d/l},m=1,...,l\},$ (8)
${\bm{z}}_{q_{i}}^{(m)}=W_{\text{out}}^{\top}\operatorname*{arg\,min}_{\bm{c}_{j}\in\mathcal{C}_{m}}||W_{\text{in}}^{\top}\bm{z}_{e_{i}}^{(m)}-\bm{c}_{j}||_{2}.$ (9)
Here $W_{\text{in}}\in\mathbb{R}^{\frac{d}{l}\times u}$, $W_{\text{out}}\in\mathbb{R}^{u\times\frac{d}{l}}$, and $u\ll d/l$. The projected vector $W_{\text{in}}^{\top}\bm{z}_{e_{i}}^{(m)}$ and codebook $\mathcal{C}_{m}$ are both $l_{2}$ normalized before finding the nearest codeword. This equalizes the scales of encoded vectors and codewords, enhancing codebook optimization by allowing a larger subset of codewords to receive gradient updates Łańcucki et al. (2020).
Pre-training paradigm. Training transformers can be challenging, and joint training with VQs is even harder. Therefore, we propose a pre-training paradigm that includes a warm-start to enhance the learning process. Initially, all VQ layers are deactivated, so no quantization occurs. Only the encoder and decoder are updated within the CS-RVQ framework, allowing latent features to bypass the quantizers and flow directly into the decoder layers. We term this the "pre-training" stage. Once the encoder and decoder have converged by minimizing reconstruction objectives, we proceed to train the entire VQ codec as usual. This approach helps mitigate the distribution shift of encoder features by optimizing the encoder in advance. It can stabilize codebook training and improve bitrate efficiency. Additionally, pre-training encoders and decoders is simpler, as it avoids the quantization errors associated with VQs. The detailed algorithm is illustrated in Appendix A.
3.6 Training Objectives
To train our codec, we employ a combination of reconstruction loss $\mathcal{L}_{recon}$ and vector quantization loss $\mathcal{L}_{vq}$. The reconstruction loss, $\mathcal{L}_{recon}$, consists of two components: an $l_{2}$ distance between the complex spectrum $\mathcal{X}$ and its reconstruction $\hat{\mathcal{X}}$, weighted by $\lambda_{1}$, and a multi-scale mel-spectrogram loss (Kumar et al., 2023), weighted by $\lambda_{2}$. These are denoted as $\mathcal{L}_{stft}$ and $\mathcal{L}_{mel}$, respectively:
$\mathcal{L}_{recon}=\lambda_{1}\mathcal{L}_{mel}+\lambda_{2}\mathcal{L}_{stft}.$ (10)
The vector quantization loss, $\mathcal{L}_{vq}$, comprises standard codebook and commitment losses as described in Eq. 2. It is averaged within the $l$ product VQs at each bitstream and summed across $s$ transmitted bitstreams. The final objective for end-to-end training is the summation of $\mathcal{L}_{recon}$ and $\mathcal{L}_{vq}$.
Figure 2: Performance evaluation results of different baseline codecs: dashed lines represent DAC baselines and solid lines represent our ESC models, with x-axis being transmission bits per second and y-axis being PESQ ()($\uparrow$)( ↑ ), Mel-Distance ()($\downarrow$)( ↓ ) and SI-SDR()($\uparrow$)( ↑ ). The metrics are averaged over our composed 1158 10-second speech clips.
4 Experiments
4.1 Experimental Setup
Datasets. We extract 150 hours 16kHz multilingual clean speech from the DNS Challenge 2021 dataset Reddy et al. (2021). Training samples are clipped into 3-second segments, and validation samples are clipped into 10-second segments. For evaluation, we compile 1158 multilingual 10-second speech clips from the LibriSpeech Panayotov et al. (2015), Multilingual LibriSpeech Pratap et al. (2020), and AIShell Shi et al. (2020) datasets. The speakers are non-overlapping in each split.
Baselines. We compare our codec against the current state-of-the-art time-domain codec DAC, by reproducing three versions [^2] Reproduction settings are detailed in Appendix B.1. on our dataset:
1) DAC-Base (Adversarial): Descript’s original released codec, operating on 16kHz audio signals. It has 74M parameter count in total.
2) DAC-Tiny (Adversarial): A smaller version of DAC-Base, with reduced encoder and decoder dimensions, for a fair comparison with ESC.
3) DAC-Tiny (Non-Adversarial): A smaller and non-adversarial version of DAC to assess the impact of discriminator on improving audio fidelity.
Implementation details. Similar to DAC baselines, we provide different versions of ESC [^3] Complete configurations are detailed in Appendix B.2.:
1) ESC-Base (Non-Adversarial): This base version consists of 6 encoder/decoder layers, with bitrates ranging from 1.5 to 9.0 kbps, and contains 8.39M parameters when operating at 9.0 kbps.
2) ESC-Base (Adversarial): Trained adversarially using the same multi-scale waveform and spectrogram discriminator as DAC.
3) ESC-Large (Non-Adversarial): A scaled-up version of ESC-Base with increased Swin Transformer layer depth, resulting in 15.58M parameters when operating at 9.0 kbps.
The ESC variants are trained using the AdamW optimizer with a learning rate of 1e-4 and betas set to (0.9, 0.999). Training runs up to 0.4 million iterations without learning rate schedulers. The proposed pre-training phase consists of 0.75 million iterations. After pre-training, the codebooks are initialized with a Kaiming normalization distribution He et al. (2015). The quantization dropout rate is set to 0.75. Loss weighting hyperparameters are set as $\lambda_{1}=0.25$, $\lambda_{2}=1.0$, and the commitment loss weighting $\beta=0.25$. In ESC-Base (Adversarial), the $\mathcal{L}_{stft}$ component is replaced with the $l_{1}$ feature matching loss, following DAC’s approach. The adjusted loss weights are $\lambda_{1}=15.0$, $\lambda_{2}=2.0$, and $\beta=0.25$.
Automatic evaluation metrics. We use objective metrics to efficiently evaluate reconstruction performance. These include the PESQ score Union (2007) from the speech enhancement domain, following Jiang et al. (2022a), the $L_{1}$ distance between log mel-spectrograms of reference and decoded waveforms (mel distance) Kumar et al. (2023), and the scale-invariant source-to-distortion ratio (SI-SDR) Le Roux et al. (2019). To measure codec inference latency, we use the real-time factor (RTF), defined by Défossez et al. (2023) as the ratio of speech audio duration to model processing time.
4.2 Comparison with DAC
We present a comprehensive comparison focusing on compression performance and inference efficiency, as demonstrated in Figure 2 and Table 1.
Performance evaluation. First, note ESC-Base and DAC-Tiny are similar in model size, each with approximately 8 million trainable parameters. Our results show that ESC-Base consistently outperforms DAC-Tiny across all bitrates, even without an adversarial discriminator. In contrast, DAC-Tiny’s reconstruction quality drops significantly without a discriminator, particularly in SI-SDR statistics. This indicates a heavy reliance of DAC models on GANs for maintaining high reconstruction quality. Notably, ESC-Base is compatible with the same convolution-based GAN discriminator used in DAC, as evidenced by its improved performance across all metric curves in its adversarial variant. Besides, ESC-Large demonstrates that increasing ESC’s model size can further enhance performance, with its PESQ curve matching that of DAC-Base, the top-performing and largest model. While DAC-Base achieves higher SI-SDR values, ESC-Large has a smaller mel-distance. Thus, we conclude that the two codecs achieve comparable performance, even though ESC-Large is trained without GANs.
Real Time Factor $\uparrow$
Codec Bitrate #Param. Enc. Dec.
ESC-Base 3.0 kbps 8.10M 33.66 34.97
6.0 kbps 8.21M 27.84 33.02
9.0 kbps 8.39M 24.45 33.95
DAC-Tiny 3.0 kbps 7.96M 42.26 49.52
6.0 kbps 8.07M 44.66 48.63
9.0 kbps 8.17M 43.00 49.10
ESC-Large 3.0 kbps 15.30M 17.91 20.81
6.0 kbps 15.41M 15.48 19.87
9.0 kbps 15.58M 13.73 20.56
DAC-Base 3.0 kbps 73.99M 12.77 3.36
6.0 kbps 74.15M 11.43 3.13
9.0 kbps 74.31M 11.81 3.25
Table 1: Complexity evaluation results of different baseline codecs: RTFs are measured from 100 10-second speech clips on an Intel Xeon Platinum 8352V CPU.
Complexity evaluation. Despite their exceptional performance, Descript’s top-performing codecs face significant computational challenges. This is evident from Table 1, where the decoding real-time factor (RTF) for DAC-Base is approximately 3.0, making it impractical for real-time applications. In contrast, our transformer-based ESC achieves much higher RTFs, indicating superior computational efficiency. Although ESC-Base is not as fast as DAC-Tiny due to the overhead of attention computation, it offers substantially better speech reconstruction capabilities, striking a favorable balance between compression performance and latency. Moreover, following the CS-RVQ scheme, ESC possesses enhanced encoding speeds at lower bitrates—a capability not found in DAC models.
These results suggest that our transformer-based codec, equipped with CS-RVQ, is a more parameter-efficient foundational codec compared to time-domain convolutional codecs. We can conclude that ESC is a more lightweight and effective neural codec, because ESC-Large achieves comparable performance to DAC-Base without requiring a powerful GAN discriminator, with approximately $\times$4.8 smaller model size, $\times$1.4 times encoding speed, and $\times$6.4 faster decoding speed.
4.3 Ablation Study
To investigate the effectiveness of the proposed components in ESC, we conducted thorough ablation experiments [^4] Implementation setups are detailed in Appendix B.3. by training frequency-domain codecs with different architectures. For fair comparisons, all other ablation models listed in Table 2 have similar model sizes to ESC-Base.
Swin Transformers and CNNs. To demonstrate that transformers are superior backbones for encoder/decoder in speech coding, we focus on two pairs of experiments: CNN/SwinT + RVQ and CNN/SwinT + CS-RVQ. In these experiments, the channel dimensions of the CNN blocks are set to match the hidden dimensions of the Swin Transformer Blocks (STBs). The comparison, as shown in the results, reveals that transformer-based codecs consistently outperform CNN-based codecs across all performance metrics and bitrates, regardless of the quantization scheme used.
CS-RVQ and RVQ. Table 2 also highlights that CS-RVQ is a superior quantization scheme compared to RVQ, regardless of whether the backbone is CNN or STB. RVQ-based codecs would hit performance bottlenecks, as adding more VQs does not improve audio quality (e.g., from 6.0kbps to 9.0kbps). However, codecs using the CS-RVQ scheme do not face such bottlenecks at higher bitrates and consistently outperform their RVQ counterparts. CS-RVQ is thus a more efficient scheme that leverages multi-scale features effectively.
Effect of pre-training paradigm. To evaluate the efficacy of a pre-training stage, we conducted an experiment of ESC-Base w/o pre-training. We monitored the VQ utilization rate, calculated as the sum of entropy (in bits) divided by the maximum number of bits from all transmitted bitstreams. This metric reflects bitrate efficiency and the fraction of seldom-used codewords. The results indicate that all experiments with pre-training achieve a near 1.0 utilization rate. However, ESC-Base w/o pre-training displays a lower utilization rate at 9.0kbps, and its reconstruction performance is also worse than ESC-Base. These findings suggest that including a pre-training stage indeed helps avoid bitrate wastage and improve audio reconstruction quality.
Method Bitrate PESQ $\uparrow$ Mel dist. $\downarrow$ SI-SDR $\uparrow$ VQ util. $\uparrow$
CNN + RVQ 3.0 kbps 2.71 2.82 0.57 96.8%
6.0 kbps 2.93 2.69 1.03 98.2%
9.0 kbps 2.96 2.68 1.05 98.7%
CNN + CS-RVQ 3.0 kbps 2.70 2.81 2.19 96.6%
6.0 kbps 3.47 2.41 3.79 97.7%
9.0 kbps 3.75 2.25 4.16 97.3%
SwinT + RVQ 3.0 kbps 2.97 2.22 0.77 98.1%
6.0 kbps 3.14 2.08 1.35 99.0%
9.0 kbps 3.16 2.07 1.39 99.2%
ESC-Base (SwinT + CS-RVQ) 3.0 kbps 3.07 2.21 3.55 97.8%
6.0 kbps 3.73 1.80 4.74 98.3%
9.0 kbps 3.92 1.62 5.33 97.9%
ESC-Base w/o Pre-training 3.0 kbps 3.09 2.25 1.75 97.7%
6.0 kbps 3.53 1.97 2.87 98.1%
9.0 kbps 3.58 1.89 2.88 86.5%
Table 2: Performance evaluation of different ablation models: results are obtained from the 1157 10-second speech clips in our test dataset.
5 Conclusions
In this paper, we introduced ESC, the first fully transformer-based neural speech codec designed for multilingual speech coding. ESC surpasses the existing state-of-the-art time-domain codecs in computational efficiency and achieves comparable compression performance even without a powerful adversarial discriminator. Our extensive evaluations demonstrate that the cross-scale residual vector quantization scheme and the Swin Transformer backbones are more efficient for neural speech coding than the traditional convolutional blocks and residual vector quantization. Overall, our study suggests a promising direction for speech compression. Future research can focus on multi-scale quantization techniques and explore additional transformer variants suitable for speech signal modeling.
6 Limitations
First, recent audio codecs are increasingly used in downstream generation tasks, where the codec acts as a foundational model to create discrete acoustic representations Borsos et al. (2023); Kreuk et al. (2022); Siuzdak (2023); Wang et al. (2023); Du et al. (2024). These compressed representations are treated as audio tokens and can be used for language modeling, making them more suitable for generation tasks. Our work does not address this significant aspect. Therefore, a promising future direction is to re-evaluate the ESC models in similar downstream applications. We believe that the cross-scale representations learned from transformer backbones would offer advantages over the fixed-scale features of convolutional codecs when applied to speech synthesis tasks. Second, different automatic metrics for audio evaluation often produce inconsistent results, which is also evidenced in our experiments. To strengthen our conclusions, it is necessary to conduct subjective evaluations involving human evaluators, such as MUSHRA listening tests Series (2014). Third, given the scalability of transformers Kaplan et al. (2020), further increasing the ESC model size and training it on more extensive audio datasets is another direction to enhance compression performance. Lastly, as discussed in previous sections, the proposed cross-scale residual vector quantization scheme requires partial decoder layers during compression, which introduces additional latency as the bitrate increases. Future research could focus on refining this scheme and developing alternative multi-scale quantization structures that are more computationally efficient.
Appendix A Pre-training Paradigm
The details of the proposed pre-training paradigm for efficiently learning VQ-based codecs are depicted in Algorithm 3. The training includes a warm start by-passing all the quantization modules and updating encoder and decoder layers only, using the reconstruction losses. In implementation, we set the quantization output $\bm{z}_{q}$ to be equal to the encoded representation $\bm{z}_{e}$. We also zero out the VQ losses to prevent codebook updates.
Algorithm 3 Pre-training Paradigm
1:repeat
2:     $\hat{\mathcal{X}}=G_{\psi}(F_{\phi}(\text{STFT}(\bm{x})))$
3:     $\mathcal{L}=\mathcal{L}_{recon}(\bm{x},\text{I-STFT}(\hat{\mathcal{X}}))$
4:     take gradient descent step on $\nabla_{\phi}\mathcal{L},\nabla_{\psi}\mathcal{L}$
5:until converged
6:activate VQs and continue learning as usual
Appendix B Experiment Details
B.1 DAC Reproduction Setups
Our customized reproduction of DAC models largely follows their released development codes. The original DAC model, designed for 16kHz audio signals, uses 12 VQs, supporting bitrates from 0.5kbps to 6.0kbps. To compare with ESC at the same bitrate levels, we extended the number of codebooks in its RVQ to 18, creating DAC-Base. For DAC-Tiny, we reduced the encoder dimension from 64 to 32 and the decoder dimension from 1536 to 288, while keeping other parameters unchanged. All these DAC baselines were trained for 0.4 million iterations with a batch size of 16 on our dataset. All other configurations can be referred to from the official release [^5] DAC’s official configuration is available here..
B.2 ESC Architecture Configurations
The configurations for ESC-Base are detailed in Table 3. The layer depth indicates the number of STBs cascaded at each encoder/decoder layer. We use GELU activation functions and LayerNorm for normalization. During quantization, our codec processes two overlapped spectrum frames together (i.e., the flattened spectrum feature is reshaped from $\mathbb{R}^{W_{i}\times H_{i}C_{i}}\rightarrow\mathbb{R}^{W_{i}/2\times 2H_{i}C_{i}}$). Overall, the vector quantizer of ESC-Base at each bitstream uses $\log_{2}1024\times 3\times 150=4500$ bits for every 3-second input speech (1.5kbps). For ESC-Large, we increase the layer depth to 4 while keeping other settings the same. In the variant ESC-Base (Adversarial), the GAN discriminator intervenes in training only after the pre-training stage ends. The GAN discriminator follows the same architecture used in DAC trainings. All three variants are trained in distributed setups for 0.4 million iterations on 4 NVIDIA RTX 4090 GPUs with a batch size of 36, taking approximately 150 GPU hours.
B.3 Details on Ablation Experiments
All ablation models were trained for 0.25 million iterations. The ablation model SwinT+CS-RVQ is exactly our ESC-Base codec. The pre-training stages constitute for 0.025 million iterations. The Swin Transformer encoder and decoder configurations follow those in ESC-Base, as demonstrated in Table 3. CS-RVQ ablation codecs follow the same configuration in Table 3 as well. For CNN-based codecs, their channel dimensions are set to match those of SwinT-based codecs. Each CNN block comprises a residual unit and a downsampling/upsampling 2D convolutional layer with a stride of 2 along the frequency dimension. The residual unit consists of two 2D convolutional layers, with a BatchNorm layer and Parametric ReLU activations interleved between them. For codecs based on RVQs, we adapted the vanilla RVQs commonly used in time-domain codecs. To handle 2-dimensional spectrum latents, we perform RVQ within different sub-channels. Specifically, the flattened feature frame is first split into sub-group vectors, similar to product VQs. Each sub-group vector is recursively quantized, as in RVQs. We set the number of product VQs to 3 and the number of residual VQs to 6, ensuring the bitrate levels match those of ESC-Base (1.5kbps per bitstream, 6 in total). Other training configurations are the same as ESC-Base. All these ablation experiments together took about 100 GPU hours on 4 RTX 4090 GPUs.
Modules Parameters Values
STFT Window/Hop Length [20ms, 5ms]
Number of FFT 382
Encoder/Decoder Patch Size [3, 2]
Layer Dims $C_{1},...,C_{6}$ [45, 72, 96, 144, 192, 384]
Attention Heads [3, 3, 6, 12, 24, 24]
Layer Depth 2
Scaling Factor $v$ 2
Vector Quantization Product VQ Size $l$ 3
Codebook Dimension $u$ 8
Codebook Size $K$ 1024
Table 3: Parameter configurations of codec variant ESC-Base, which comprises 6 encoder/decoder layers.

### v3

@@
 Neural speech codecs aim to compress input signals into minimal bits while maintaining content quality in a low-latency manner.
-However, existing codecs often trade computational complexity for reconstruction performance.
-These codecs primarily use convolutional blocks for feature transformation layers, which are not inherently suited for capturing the local redundancies in speech signals.
+However, existing neural codecs often trade model complexity for reconstruction performance.
+These codecs primarily use convolutional blocks for feature transformation, which are not inherently suited for capturing the local redundancies in speech signals.
 To compensate, they require either adversarial discriminators or a large number of model parameters to enhance audio quality.
-In response to these challenges, we introduce the Efficient Speech Codec (ESC), a lightweight, parameter-efficient speech codec based on a cross-scale residual vector quantization scheme and transformers.
+In response to these challenges, we introduce the Efficient Speech Codec (ESC) [^1] Code and pretrained models available at https://github.com/yzGuu830/efficient-speech-codec, a lightweight, parameter-efficient speech codec based on a cross-scale residual vector quantization scheme and transformers.
 Our model employs mirrored hierarchical window transformer blocks and performs step-wise decoding from coarse-to-fine feature representations.
 To enhance bitrate efficiency, we propose a novel combination of vector quantization techniques along with a pre-training paradigm.
-Extensive experiments demonstrate that ESC can achieve high-fidelity speech reconstruction with significantly lower complexity, making it a promising alternative to existing convolutional audio codecs [^1] Our code and pretrained models are available here..
+Extensive experiments demonstrate that ESC can achieve high-fidelity speech reconstruction with significantly lower model complexity, making it a promising alternative to existing convolutional audio codecs.
 ESC: Efficient Speech Coding with Cross-Scale Residual Vector Quantized Transformers
@@
 (2015).
-Early efforts integrating deep generative models such as WaveNet Oord et al.
+Early efforts integrating deep generative models, such as WaveNet Oord et al.
 (2016) and SampleRNN Mehri et al.
-(2017) into audio codecs have yielded promising results.
-These models serve as powerful decoders, synthesizing high-quality speech audio based on the intermediate representations generated by traditional codecs Kleijn et al.
+(2017), into audio codecs have delivered promising results.
+These models, acting as powerful decoders, synthesize high-quality speech from intermediate representations produced by traditional codecs Kleijn et al.
 (2018); Klejsa et al.
 (2019).
-However, their auto-regressive decoding nature often introduces significant inference latency, limiting their practical application.
-Alternatively, some end-to-end neural audio codecs build upon the vector quantization (VQ) network first introduced by Van Den Oord et al.
+However, their auto-regressive nature of the decoding process often introduces significant inference latency, limiting their practical application.
+Alternatively, some end-to-end neural audio codecs leverage the vector quantization (VQ) network first introduced by Van Den Oord et al.
 (2017).
-VQ networks, which consist of trainable codebooks, can discretize continuous input vectors by assigning them to the nearest neighboring code-vectors, making them suitable for both generation and compression tasks.
-Based on this framework, existing VQ codecs Zeghidour et al.
+VQ networks use a learnable collection of code-vectors, known as a codebook, to quantize continuous vectors by assigning them to the nearest codeword.
+This discretization positions VQNs well-suited for both generation and compression tasks.
+Following this approach, existing VQ codecs Zeghidour et al.
 (2021); Défossez et al.
 (2023); Kumar et al.
-(2023) typically follow a three-stage architecture with a convolutional encoder and decoder interleaved with a residual vector quantization (RVQ) module Vasuki and Vanathi (2006) at the bottleneck.
-The encoder and decoder perform downsampling and upsampling of audio signals in the time-domain to produce hierarchical representations.
-RVQ is an enhanced quantization scheme where a series of VQs recursively quantize the residuals from previous stages.
-Additionally, these codecs utilize generative adversarial networks (GANs) to remove artifacts and produce high-fidelity audio reconstructions.
-Significant effort has been dedicated to developing effective audio discriminators that leverage multi-scale features of waveforms and spectrograms to improve the generated audio during training Esser et al.
-(2021); Kumar et al.
-(2019); Kong et al.
+(2023) typically employ a three-stage architecture: a convolutional encoder and decoder, and a residual vector quantization (RVQ) module Vasuki and Vanathi (2006) applied in the latent space.
+The encoder and decoder downsample and upsample audio waveform features, creating hierarchical representations.
+RVQ further refines vanilla vector quantization by minimizing quantization error through a series of codebooks that recursively quantize the residuals from previous stages.
+Additionally, these codecs employ adversarial discriminators to remove artifacts and produce high-fidelity audio reconstructions.
+Substantial effort has been dedicated to designing effective audio discriminators, including an improved feature matching loss Kumar et al.
+(2019), as well as various multi-resolution waveform and spectrogram discriminators Kong et al.
+(2020); Zeghidour et al.
+(2021); Défossez et al.
+(2023); gil Lee et al.
 (2023).
-Currently, VQ-based codecs have demonstrated remarkable capabilities in audio reconstruction even at ultra-low bitrates.
-Despite these advantages, we find that convolutional VQ-based codecs heavily rely on powerful GAN discriminators to produce high-quality audio, introducing additional optimization challenges due to adversarial training.
-Moreover, convolutional codecs tend to face computational limitations, as they require a large number of parameters to achieve a good trade-off between compression rate and reconstruction quality.
-To address these issues, our work aims to develop a more parameter-efficient speech codec by implementing the following architectural improvements: 1) Replacing convolutional layers with efficient Swin-Transformer Blocks (STBs) Liu et al.
-(2021), which have greater capacity for modeling acoustic features; 2) Utilizing the cross-scale residual vector quantization (CS-RVQ) scheme Jiang et al.
+VQ-based audio codecs have demonstrated remarkable performance in audio reconstruction, even at ultra-low bitrates.
+Despite these advantages, we find that convolutional VQ codecs heavily depend on powerful discriminators to produce high-quality audio, posing additional optimization challenges due to adversarial training.
+Moreover, these codecs tend to confront computational constraints, as they require a large number of parameters to balance high compression rates and reconstruction performance.
+To address these issues, our work develops a more parameter-efficient speech codec by reducing model complexity and implementing the following architectural improvements: 1) replacing convolutional layers with efficient Swin-Transformer Blocks (STBs) Liu et al.
+(2021), which can better model acoustic features; 2) utilizing the cross-scale residual vector quantization (CS-RVQ) scheme Jiang et al.
 (2022a) instead of RVQ, extending quantization from a fixed level to multiple levels.
-In addition, training VQ-based codecs encounters a significant obstacle: a fraction of the vector quantizer’s codebook is often underutilized for representing input vectors, leading to a phenomenon known as codebook collapse.
-This issue is frequently observed in image generation tasks Takida et al.
+In addition, training VQ codecs frequently leads to a significant challenge: codebook collapse, where a fraction of the codebook remains underutilized in representing input vectors.
+This issue is frequently observed when training visual tokenizers for generative vision tasks Takida et al.
 (2022); Zhang et al.
@@
 (2023).
-To address this in the context of speech compression, we propose combining product vector quantization (PVQ) Baevski et al.
-(2019), latent dimension reduction Yu et al.
+To address this problem in speech compression, we propose combining product vector quantization (PVQ) Baevski et al.
+(2019), code factorization Yu et al.
 (2022), and Euclidean normalization Łańcucki et al.
 (2020) to enhance codebook utilization.
-Additionally, we introduce a learning paradigm to facilate codebook optimization.
-It incorporates a pre-training stage during which VQs are freezed and trained afterwards.
-In summary, the contributions of our work are as follows:
-  • We introduce ESC, a lightweight efficient speech codec that integrates cross-scale residual vector quantization (CS-RVQ) and mirrored hierarchical Swin-Transformer blocks, achieving a superior trade-off between reconstruction performance and codec complexity compared to current state-of-the-art models.
-  • We propose a novel combination of vector quantization techniques along with a pre-training paradigm that includes a warm start for the encoder and decoder, aimed at mitigating codebook collapse and enhancing bitrate efficiency.
-  • Our extensive experiments demonstrate that transformers and CS-RVQ, the core components of ESC, are more efficient backbones for speech compression than convolutions and RVQ.
+Furthermore, we introduce a learning paradigm to facilitate optimization, which includes a pre-training stage where the codebooks are deactivated and trained subsequently.
+In summary, the key contributions of our work are as follows:
+  • We introduce ESC, a fully transformer-based speech codec with cross-scale quantization structures.
+It achieves a superior tradeoff between compression rate, reconstruction quality, and model complexity, outperforming current state-of-the-art models.
+  • We propose a novel combination of vector quantization techniques within the cross-scale residual vector quantization (CS-RVQ) framework, coupled with a pre-training paradigm that effectively mitigates codebook collapse and enhances bitrate efficiency.
+  • Extensive comparisons with Descript’s audio codec on a multilingual speech corpus demonstrate that transformers and CS-RVQ, the core components of ESC, are superior backbones for speech foundation models than the mainstream convolutions and RVQ.
 2 Related Work
-2.1 VQ-based Audio Codecs
-Notable examples in recent VQ-based audio codecs include SoundStream Zeghidour et al.
+2.1 Neural Audio Codecs
+Recently, most notable neural audio codecs have been based on the vector quantization (VQ) network, including SoundStream Zeghidour et al.
 (2021), EnCodec Défossez et al.
@@
 (2023).
-SoundStream is recognized as the first universal neural codec designed for various audio types.
-EnCodec enhances compression rates by integrating a lightweight transformer language model at the discrete latent space and explores a streaming architecture.
-Following similar methodologies, Kumar et al.
-(2023) identify the side effects of quantization dropout, a technique for bitrate scalability, and demonstrate the superiority of a periodic inductive bias function over other common activation functions in audio signal modeling gil Lee et al.
+SoundStream is distinguished as the first universal codec capable of handling diverse audio types.
+EnCodec improves compression rates by integrating a lightweight transformer language model within the discrete latent space and implements a streaming architecture.
+Building on similar backbones, Kumar et al.
+(2023) further explore the implications of quantization dropout, a technique for bitrate scalability, and demonstrate the superiority of periodic inductive bias functions over common activation functions for audio signal modeling gil Lee et al.
 (2023); Ziyin et al.
 (2020).
-These codecs process audio signals directly in the time domain and are thus categorized as time-domain codecs.
-Meanwhile, frequency-domain codecs, such as TFNet Jiang et al.
-(2022b) and our proposed ESC, leverage the invertible short-time Fourier transform (STFT) to convert time-domain audio into spectrograms.
-The encoder and decoder thus process these two-dimensional spectral features directly.
-Among recent neural codecs, DAC achieves state-of-the-art compression ratios and audio reconstruction quality, although its computation bottleneck is ignored.
+These models directly process audio waveforms and are classified as time-domain codecs.
+In contrast, frequency-domain codecs focus on processing more intuitive audio spectrogram features.
+Lyra Kleijn et al.
+(2021), for example, converts audio waveforms into log mel-spectrograms and directly quantizes them into tokens.
+Due to the non-invertibility of mel-spectrograms, it relies on a vocoder Kalchbrenner et al.
+(2018) for waveform synthesis.
+To circumvent the inefficiencies associated with heavy vocoders, some frequency-domain codecs, including TFNet Jiang et al.
+(2022b) and our ESC, employ the invertible Short-time Fourier Transform (STFT) to convert waveforms into complex spectra.
+This design enables the reconstructed STFT spectra to be seamlessly inverted back into waveforms without information loss using inverse-STFT.
+Among recent audio codecs, DAC achieves state-of-the-art compression ratios and reconstruction quality, though its computation bottlenecks are sometimes overlooked.
 2.2 Swin Transformers
 Vision Transformers (ViTs) Dosovitskiy et al.
-(2020) have outperformed convolutional neural networks (CNNs) in various image processing tasks due to their superior ability to capture complex patterns.
+(2020) have outperformed convolutional neural networks (CNNs) in various image processing tasks, largely due to their superior ability to capture complex patterns.
 The Swin Transformer Liu et al.
-(2021), a notable variant, enhances this capability through a hierarchical approach that leverages window attention mechanisms.
-It can scale efficiently to high-resolution signals while maintaining computational efficiency.
-In image compression, Swin Transformers have demonstrated exceptional performance.
+(2021), a notable variant, enhances this capability by employing a hierarchical approach with shifted window attention mechanisms, enabling it to scale efficiently to high-resolution signals while maintaining computational efficiency.
+In the context of image compression, Swin Transformers have demonstrated exceptional performance.
 Studies by Zhu et al.
 (2021) and Zou et al.
-(2022) have shown that they outperform CNNs in modeling spatial hierarchies and long-range dependencies.
-The attention mechanisms enable the accurate preservation of essential details and textures, even at lower bitrates.
-These capabilities highlight the potential of transformers for applications beyond image compression, suggesting their suitability for the field of speech compression as well.
+(2022) show that Swin Transformers surpass CNNs in modeling spatial hierarchies and long-range dependencies.
+The attention mechanism facilitates the accurate preservation of essential details and textures, even at lower bitrates.
+These capabilities suggest that transformers could also be effective in applications beyond image compression, such as modeling audio spectrograms.
 2.3 Vector Quantization
-Proposed in Vector Quantized Variational Autoencoders (VQ-VAEs) (Van Den Oord et al., 2017), vector quantization (VQ) serves as a trainable layer that deterministically quantizes the encoded latent variables to their nearest neighbors in an embedding codebook.
-A vector quantization layer $Q(\cdot\ ;\mathcal{C})$ is parameterized by a codebook collection of continuous vectors $\mathcal{C}={\bm{c}_{1},...,\bm{c}_{K}}$.
-Each vector is referred to as a codeword, and its index is known as a code.
-The layer quantizes a vector $\bm{z}_{e}\in\mathbb{R}^{d}$ to be $\bm{z}_{q}\in\mathbb{R}^{d}$ by selecting the Euclidean nearest codeword $\bm{c}_{k}$ from the codebook $\mathcal{C}$, i.e.,
+In the Vector Quantized Variational Autoencoder (VQ-VAE) (Van Den Oord et al., 2017), vector quantization (VQ) functions as a trainable layer that deterministically quantizes encoded latent variables by mapping them to their nearest neighbors in an embedding codebook.
+A VQ layer, denoted as $Q(\cdot\ ;\mathcal{C})$, is parameterized by a collection of continuous vectors $\mathcal{C}=\{\bm{c}_{1},...,\bm{c}_{K}\}$, each referred to as a codeword, with its associated index known as a code.
+The layer quantizes a vector $\bm{z}_{e}\in\mathbb{R}^{d}$ to $\bm{z}_{q}\in\mathbb{R}^{d}$ by selecting the Euclidean nearest codeword $\bm{c}_{k}$ from the codebook $\mathcal{C}$, i.e.,
 $\bm{z}_{q}:=\bm{c}_{k}=\operatorname*{arg\,min}_{\bm{c}_{j}\in\mathcal{C}}||\bm{z}_{e}-\bm{c}_{j}||_{2}^{2}.$ (1)
-For notation convenience, we denote the vector quantization function as $\bm{z}_{e}\xrightarrow[]{Q(\cdot;\mathcal{C})}(\bm{z}_{q},\tilde{z}_{q})$, where $\tilde{z}_{q}:=k$ is the corresponding discrete code for the nearest codeword.
+For convenience, we denote the output of the VQ function as $(\bm{z}_{q},\tilde{z}_{q})$, where $\tilde{z}_{q}:=k$ represents the discrete code corresponding to the nearest codeword.
 During compression, the encoding process outputs the discrete index $\tilde{z}_{q}$, which is stored with a $\log_{2}K$ bit budget.
-The decoding process starts by retrieving the continuous vector $\bm{z}_{q}$ from the codebook using the encoded index $\tilde{z}_{q}$.
+The decoding process starts by retrieving the continuous vector $\bm{z}_{q}$ from the codebook using the index $\tilde{z}_{q}$.
 The VQ function $Q(\cdot;\mathcal{C})$ is non-differentiable due to the $\operatorname*{arg\,min}$ operator.
@@
 (2013) to bypass this in back-propagation.
-In other words, the gradient component $\frac{\partial\bm{z}_{q}}{\partial\bm{z}e}$ is eliminated, and as compensation, a codebook loss and a commitment loss are proposed to pull together the codewords and the encoded latents:
+In other words, the gradient component $\frac{\partial\bm{z}_{q}}{\partial\bm{z}e}$ is estimated by identity.
+Additionally, auxiliary losses including codebook loss and commitment loss are proposed to pull the codewords and latent features closer:
 $\mathcal{L}_{vq}=||\text{sg}(\bm{z}_{e})-\bm{z}_{q}||_{2}^{2}+\beta||\bm{z}_{e}-\text{sg}(\bm{z}_{q})||_{2}^{2}.$ (2)
-Here $\text{sg}(\cdot)$ denotes the stop gradient operator.
-The first term updates the codebook with an $l_{2}$ error, moving the codewords towards the encoder outputs.
-The second term ensures $\bm{z}_{e}$ can commit to the embedding without growing arbitrarily.
-The scalar $\beta$ is used to balance the importance of updating the codebook and encoder.
+Here $\text{sg}(\cdot)$ denotes the stop-gradient operator.
+The first term updates the codebook with an $l_{2}$ error, pushing the codewords towards the input vectors.
+The second term ensures that $\bm{z}_{e}$ commits to the embedding without growing arbitrarily.
+The scalar $\beta$ balances the importance of updating the codebook and the encoder.
 2.4 Codebook Collapse
-Straight-through estimators (STEs) often lead to significant issues, most notably the issue of codebook collapse Vuong et al.
+Straight-through estimators (STEs) can lead to significant issues, most notably codebook collapse, as detailed by Vuong et al.
 (2023).
 In a recent study, Huh et al.
-(2023) offer a straightforward explanation, attributing it to an internal codebook covariate shift during the training phase.
-Frequent adjustments in encoder representations cause misalignment with VQ codebooks, resulting in only a subset of codewords being updated.
-Consequently, VQ layers are prone to divergence, often ending up with a significant number of dead vectors.
-Various strategies have been proposed in generative modeling tasks to address this problem, including the use of stochastic VQs Takida et al.
+(2023) provide a plausible explanation, attributing the collapse to an internal codebook covariate shift during training.
+Frequent adjustments in encoder representations cause misalignment with the codebook, resulting in only a subset of codewords being updated.
+Consequently, VQ layers are prone to divergence, often ending up with a significant number of inactive vectors.
+Various strategies have been proposed in generative modeling context to address this issue, including stochastic quantization Takida et al.
 (2022); Zhang et al.
-(2023), self-annealing soft-to-hard quantization approaches Agustsson et al.
-(2017), re-initializing codewords with K-means centroids every few iterations Łańcucki et al.
+(2023), self-annealed soft-to-hard quantization Agustsson et al.
+(2017), re-initializing codewords using K-means centroids every few epochs Łańcucki et al.
 (2020); Dhariwal et al.
-(2020), and avoiding the issues of STEs with finite scalar quantizers Mentzer et al.
+(2020), and reformulating with finite scalar quantization Mentzer et al.
 (2024).
 In audio compression, Kumar et al.
-(2023) attempt to mitigate codebook collapse by reducing the dimensions of codewords Yu et al.
+(2023) address codebook collapse by down-projecting codewords Yu et al.
 (2022) and normalizing them within a Euclidean ball Łańcucki et al.
@@
 Encoder STBs iteratively halve the frequency resolution and produce hierarchical feature representations.
-Mirrored decoder STBs iteratively recover the frequency resolution by leveraging coarse-to-fine encoded features.
-The intermediate residuals between encoder and decoder hidden states are progressively quantized to refine decoding.
-The entire network is solely composed of efficient transformer blocks.
+Mirrored decoder STBs recover the frequency resolution by progressively leveraging coarse-to-fine quantized residual features between encoder and decoder hidden states.
+The entire network is solely composed of efficient transformer blocks and vector quantization layers.
 The figure displays a scenario when the deepest $3$ of $n+1$ total bitstreams (solid lines) are transmitted, with others left inactive.
@@
 3.1 Overall Architecture
-As illustrated in Figure 1, our codec operates on the complex spectrum $\mathcal{X}\in\mathbb{R}^{2\times F\times T}$ derived from the Short-Time Fourier Transform (STFT) of a speech signal.
+As illustrated in Figure 1, ESC operates on the complex spectrum $\mathcal{X}\in\mathbb{R}^{2\times F\times T}$ derived from the Short-Time Fourier Transform (STFT) of a speech signal.
 Here, the real and imaginary components of $\mathcal{X}$ are treated as separate channels.
-The model consists of a series of mirrored transformer encoder and decoder layers, each performing downsampling or upsampling to generate coarse and fine representations, as detailed in Section 3.3.
-Starting from the quantized latents at the bottleneck VQ, the decoder progressively reconstructs the original spectrum by leveraging multi-level residuals between the intermediate features of the encoder and decoder.
-This cross-scale decoding mechanism is explained in further detail in Section 3.4.
-Finally, the reconstructed spectrum $\hat{\mathcal{X}}$ is transformed back into the time domain through the inverse STFT.
+Instead of using strided convolutions, ESC comprises a series of mirrored transformer encoder and decoder layers, each performing downsampling or upsampling to create coarse and fine representations, as described in Section 3.3.
+Starting from the quantized latents at the bottleneck VQ, the decoder progressively reconstructs the original spectrum by leveraging multi-level quantized residuals between the intermediate features of the encoder and decoder.
+This cross-scale decoding mechanism is further detailed in Section 3.4.
+Finally, the reconstructed spectrum $\hat{\mathcal{X}}$ is transformed back into a waveform through the inverse-STFT.
 3.2 Notations
-We first define some architectural notations for clarity.
-We represent the encoder and decoder functions $F_{\phi}(\cdot)$ and $G_{\psi}(\cdot)$, each as a composition of individual layer functions $f_{\phi_{1}},\ldots,f_{\phi_{n}}$ and $g_{\psi_{1}},\ldots,g_{\psi_{n}}$, respectively.
-We use $\mathcal{Z}$ to represent spectrum feature and $\bm{z}$ to denote a feature frame.
-Specifically, we denote $\mathcal{Z}_{e_{i}}$ as the hidden feature after $i$ encoder layers, and $\mathcal{Z}_{q_{i}}$ as the corresponding $i$-th decoder feature.
-Formally, this is expressed as:
+We first define some notations for clarity.
+The encoder and decoder are denoted by $F_{\phi}(\cdot)$ and $G_{\psi}(\cdot)$, respectively, each being a composition of individual layer functions $f_{\phi_{1}},\ldots,f_{\phi_{n}}$ and $g_{\psi_{1}},\ldots,g_{\psi_{n}}$.
+We use $\mathcal{Z}\in\mathbb{R}^{C\times F\times T}$ to denote a spectrum feature and $\bm{z}\in\mathbb{R}^{CF}$ to denote a flattened time frame vector in $\mathcal{Z}$.
+Specifically, $\mathcal{Z}_{e_{i}}$ refers to the feature after the $i$-th encoder layer, and $\mathcal{Z}_{q_{i}}$ denotes the $i$-th decoder feature.
 $\mathcal{Z}_{e_{i}}=f_{\phi_{i}}\circ...\circ f_{\phi_{1}}\circ\mathcal{Z}_{e_{0}}$ (3)
 $\mathcal{Z}_{q_{i}}=g_{\psi_{i}}\circ...\circ g_{\psi_{1}}\circ\mathcal{Z}_{q_{0}},$ (4)
-where $\mathcal{Z}_{e_{0}}$ is the original input feature and $\mathcal{Z}_{q_{0}}$ is the latent representation at the bottleneck.
+Here, $\mathcal{Z}_{e_{0}}$ is the original input feature and $\mathcal{Z}_{q_{0}}$ is the latent representation at the bottleneck.
 3.3 Transformer Encoder and Decoder
 To effectively capture redundancies within audio signals, we replace convolutional layers with hierarchical Swin Transformer blocks (STBs) and their extended decoding counterparts.
-Patchify. The encoder begins with a linear patchify module, where the complex spectrum $\mathcal{X}$ is divided into small patches and linearly projected:
+Patchify. The encoder starts with a linear patchify module, where the complex spectrum $\mathcal{X}$ is divided into small patches and linearly up-projected:
 […]
 Here, the patch size across the frequency and temporal dimensions is $(\frac{F}{H_{0}},\frac{T}{W_{0}})$.
-This step reduces the input resolution to alleviate the computational burden on transformers.
-At the end of the decoder, a symmetric de-patchify module reshapes the decoded patch embedding $\mathcal{Z}_{q_{n}}$ and linearly projects it back to produce a recovered spectrum $\hat{\mathcal{X}}$.
+This step reduces the input resolution to alleviate the computational burden on attention computation.
+At the end of the decoder, a symmetric de-patchify module reshapes the decoded patch feature $\mathcal{Z}_{q_{n}}$ and linearly down-projects it to produce a recovered spectrum $\hat{\mathcal{X}}$.
 Swin Transformer blocks.
@@
 Downsampling and upsampling.
-Our codec maintains temporal resolution while scaling frequency resolution to equalize bitrates across different bitstreams.
+ESC maintains temporal resolution while scaling frequency resolution to equalize bitrates across different bitstreams.
 To achieve this, we modify the original patch merging/splitting modules with a single-dimensional pixel unshuffle/shuffle module Shi et al.
 (2016) along the frequency dimension.
-During encoder downsampling, an intermediate encoder spectrum feature $\mathcal{Z}_{e_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is first reshaped and then projected as follows:
+During encoder downsampling, an intermediate encoder spectrum feature $\mathcal{Z}_{e_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is first reshaped and then projected by $P_{e_{i}}\in\mathbb{R}^{vC_{i}\times C_{i+1}}$ as follows:
 […]
-where $v$ is the scaling factor.
+where $v$ is the down-scaling factor.
 The upsampling process mirrors this operation in reverse.
-An intermediate decoder feature $\mathcal{Z}_{q_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is up-projected and then reshaped, resulting in an up-scaled frequency resolution:
+An intermediate decoder feature $\mathcal{Z}_{q_{i}}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ is first projected by $P_{q_{i}}\in\mathbb{R}^{C_{i}\times vC_{i+1}}$ and then reshaped, resulting in an up-scaled frequency resolution:
 […]
-Thus, the transformer encoder and decoder layers are mirrored, creating symmetric and hierarchical representations of the input audio spectrum.
+Overall, the transformer encoder and decoder layers are mirrored, creating symmetric and hierarchical representations of the input audio spectrum.
 With these backbones, ESC is a fully transformer-based codec without any convolutional modules.
 3.4 Cross-Scale Residual Vector Quantization
-For efficient modeling of audio signals, ESC employs multi-scale features through a coarse-to-fine decoding process to generate bitstreams at different levels.
-Our codec integrates the more intuitive cross-scale residual vector quantization (CS-RVQ) structure proposed by Jiang et al.
-(2022a), eliminating the need for additional networks to fuse encoder and decoder features.
-As depicted in Algorithm 1, Algorithm 2 and Figure 1, the decoding process is conditioned on encoder features from various resolution scales.
-This distinguishes it from the commonly used residual vector quantization scheme that operates only at the bottom scale, which often overlooks low-level information.
+To achieve parameter-efficient modeling of audio signals, ESC employs multi-scale features that capture coarse-to-fine information.
+It integrates the more intuitive residual-based cross-scale vector quantization (CS-RVQ) framework proposed by Jiang et al.
+(2022a), eliminating the need for additional networks to merge encoder and decoder features for improved reconstruction quality.
+As depicted in Algorithm 1, Algorithm 2 and Figure 1, the decoding process is conditioned on the encoded quantized residuals between encoder and decoder features from low-to-high resolution scales.
+This approach differs from the commonly used residual vector quantization scheme, which operates solely at the lowest scale, relying on the highest-level information while overlooking low-level details.
 Encoding.
-The encoding process begins with the encoder function $F_{\phi}(\cdot)$, creating multi-scale encoder features $\bm{z}_{e_{1}},...,\bm{z}_{e_{n}}$.
+The encoding process begins with the encoder $F_{\phi}(\cdot)$, creating multi-scale encoder features $\bm{z}_{e_{1}},...,\bm{z}_{e_{n}}$.
 $\bm{z}_{e_{n}}$ is first quantized by the bottleneck quantizer $Q_{0}$ to form the lowest bitstream.
-This represents the simplest case when $s$ is set to 1, and CS-RVQ reduces to a fixed-scale VQ at the bottleneck.
-At higher bitstreams, the residual between symmetric encoder and decoder features with higher resolutions, $\bm{z}_{e_{n-i+1}}-\bm{z}_{q_{i-1}}$, is quantized by $Q_{i}$.
-This quantized residual $\bm{q}_{i}$ is added back to $\bm{z}_{q_{i-1}}$ and decoded by the subsequent decoder layer $g_{\phi_{i}}(\cdot)$, producing the next decoder feature $\bm{z}_{q_{i}}$.
-Recursively, residuals at higher resolutions are progressively quantized, forming the remaining bitstreams.
-This mechanism enables multi-scale learning and allows the decoder layers to incrementally reduce quantization errors by conditioning on encoder features.
-When $s>2$, the encoding process of CS-RVQ requires $s-2$ additional decoder layers to produce residuals at higher levels.
-After encoding, $\bm{z}_{e_{0}}$ is compressed to be multi-level codes $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},\ldots,\tilde{z}_{q_{s-1}}$.
-Decoding.
-The decoding process starts by retrieving the quantized latent from the bottom VQ using the code index $\tilde{z}_{q_{0}}$, which provides the original decoder input $\bm{z}_{q_{0}}$.
-At higher levels, $\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$ are iteratively used to retrieve codewords, resulting in multi-scale quantized residuals $\bm{q}_{1},...,\bm{q}_{s-1}$.
-Each quantized residual $\bm{q}_{i}$ is added back to the corresponding decoder feature $\bm{z}_{q_{i-1}}$ to improve the decoding process.
-Starting from the $s$-th decoder layer, there are no quantized residuals, and the layers perform vanilla decoding.
-Finally, the recovered frame spectrum $\bm{z}_{q_{n}}$ is obtained, which conditions on $s-1$ residual features.
-Training.
-During the training of CS-RVQ codecs, the encoding and decoding processes are combined to form a complete forward pass.
-To enable bitrate scalability, we sample the number of bitstreams $s\sim\text{Uniform}\{1,\ldots,n\}$ at a rate $p$ within each training batch.
-$p$ is a hyperparameter that balances the reconstruction quality at different bitrates, as proposed by Kumar et al.
-(2023).
+This represents the simplest case when the number of transmitted bitstream $s$ is set to 1, and CS-RVQ reduces to a fixed-scale VQ at the bottleneck.
+For higher bitstreams, the residual between symmetric encoder and decoder at higher resolutions, $\bm{z}_{e_{n-i+1}}-\bm{z}_{q_{i-1}}$, is quantized by $Q_{i}$.
+The quantized residual $\bm{q}_{i}$ is then added back to $\bm{z}_{q_{i-1}}$ and decoded by the subsequent decoder layer $g_{\phi_{i}}(\cdot)$, producing the next decoder feature $\bm{z}_{q_{i}}$.
+Recursively, residuals at higher resolutions are progressively quantized, forming the remaining bitstreams (see Algorithm 1, Lines 3-6).
+This mechanism enables multi-scale learning, allowing the decoder layers to incrementally reduce quantization errors by conditioning on encoder-decoder residual features.
+When $s>2$, this encoding process requires forward passing $s-2$ additional decoder layers to produce residuals at higher levels.
+After encoding, the input $\bm{z}_{e_{0}}$ is compressed into multi-level codes $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},\ldots,\tilde{z}_{q_{s-1}}$.
 Algorithm 1 CS-RVQ Encoding
-1:Patch embedding frame $\bm{z}_{e_{0}}\in\mathbb{R}^{C_{0}\times H_{0}}$, encoder $F_{\phi}(\cdot)$, decoder $G_{\psi}(\cdot)$, vector quantizers $Q_{0},Q_{1},...,Q_{s-1}$, number of bitstreams $s$
+1:A flattened time frame $\bm{z}_{e_{0}}\in\mathbb{R}^{C_{0}H_{0}}$, encoder $F_{\phi}(\cdot)$, decoder $G_{\psi}(\cdot)$, vector quantizers $Q_{0},Q_{1},...,Q_{s-1}$, number of bitstreams $s$
 2:$\bm{z}_{e_{1}},...,\bm{z}_{e_{n}}\leftarrow F_{\phi}(\bm{z}_{e_{0}})$triangleright Encoder forward pass
-3:$\bm{z}_{q_{0}},\tilde{z}_{q_{0}}\leftarrow Q_{0}(\bm{z}_{e_{n}})$triangleright Bottleneck VQ
+3:$\bm{z}_{q_{0}},\tilde{z}_{q_{0}}\leftarrow Q_{0}(\bm{z}_{e_{n}})$triangleright bottom VQ
 4:for $i=1\dots s-2$ do
@@
 6:     $\bm{z}_{{q}_{i}}\leftarrow g_{\psi_{i}}(\bm{z}_{{q}_{i-1}}+\bm{q}_{i})$
-7:end $\triangleright$ Encoding requires $s-2$ decoder layers
+7:end $\triangleright$ Encoding involves $s-2$ decoder layers
 8:if $s>1$ then
@@
 1:Codes $\tilde{z}_{q_{0}},\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$, decoder $G_{\psi}(\cdot)$, vector quantizers $Q_{0},Q_{1},...,Q_{s-1}$
-2:$\bm{z}_{q_{0}}\xleftarrow[]{Q_{0}}\tilde{z}_{q_{0}}$triangleright Retrieve codewords from VQ
+2:$\bm{z}_{q_{0}}\xleftarrow[]{Q_{0}}\tilde{z}_{q_{0}}$triangleright Retrieve codewords from bottom VQ
 3:for $i=1\dots s-1$ do
@@
 5:     $\bm{z}_{{q}_{i}}\leftarrow g_{\psi_{i}}(\bm{z}_{{q}_{i-1}}+\bm{q}_{i})$
-6:end $\triangleright$ Decoding conditions on residuals
+6:end $\triangleright$ Decoding refined by quantized residuals
 7:for $i=s\dots n$ do
 8:     $\bm{z}_{{q}_{i}}\leftarrow g_{\psi_{i}}(\bm{z}_{{q}_{i-1}})$
-9:end $\triangleright$ Continue with vanilla decoding
+9:end $\triangleright$ Continue with regular decoding
 10:return $\bm{z}_{q_{n}}$
+Decoding.
+The decoding process starts by retrieving the quantized latent at the bottom VQ using code $\tilde{z}_{q_{0}}$, which provides the initial decoder input $\bm{z}_{q_{0}}$.
+At higher levels, the codes $\tilde{z}_{q_{1}},...,\tilde{z}_{q_{s-1}}$ are iteratively used to retrieve codewords, producing multi-scale low-to-high quantized residuals $\bm{q}_{1},...,\bm{q}_{s-1}$.
+In Algorithm 2, Lines 2-5, each quantized residual $\bm{q}_{i}$ is added back to the corresponding decoder feature $\bm{z}_{q_{i-1}}$ to refine the decoding process.
+Starting from the $s$-th decoder layer, there are no quantized residuals, and the remaining layers perform regular decoding.
+Finally, the recovered frame vector $\bm{z}_{q_{n}}$ is obtained, benefiting from $s-1$ quantized residual features.
+Training.
+During training, the encoding and decoding processes are concatenated to form a complete forward pass.
+To enable bitrate scalability, we sample $s\sim\text{Uniform}\{1,\ldots,n\}$ at a rate $p$ within each training mini-batch.
+$p$ is a hyperparameter that balances the reconstruction quality at different bitrates, as proposed by Kumar et al.
+(2023).
 3.5 Mitigating Codebook Collapse
-During quantization, the spectrum frame feature needs to be flattened, merging the frequency and channel dimensions.
-This results in large hidden dimensions for VQ, increasing the possibility of codebook collapse during training.
+ESC performs a per-frame vector quantization.
+Before nearest neighbor searching, each input spectrum frame feature in $\mathcal{Z}$ needs to be flattened, merging the frequency and channel dimensions.
+This approach can result in large input vector dimensions for VQ, increasing the optimization challenges associated with codebook underutilization.
 Vector quantization setups.
-To effectively optimize the codebooks, we modify vanilla VQs using a combination of product vector quantization and vector dimension reduction at each bitstream.
-Specifically, a frame feature $\bm{z}_{e_{i}}\in\mathbb{R}^{C_{i}\times H_{i}}$ is flattened into a $d$-dimensional vector before quantization.
-It is then split into a collection of $l$ sub-vectors.
-Each sub-vector $\bm{z}_{e_{i}}^{(m)}$ is down-projected and quantized by an individual codebook $\mathcal{C}_{m}$ to be ${\bm{z}}_{q_{i}}^{(m)}$:
-$\bm{z}_{e_{i}}\equiv\{\bm{z}_{e_{i}}^{(m)}\mid\bm{z}_{e_{i}}^{(m)}\in\mathbb{R}^{d/l},m=1,...,l\},$ (8)
+To optimize the codebooks effectively, we modify the vanilla VQ by combining product vector quantization with code-vector factorization at each bitstream.
+Specifically, a flattened $d$-dimensional frame vector $\bm{z}_{e_{i}}$ is split into a set of $l$ sub-vectors.
+Each sub-vector $\bm{z}_{e_{i}}^{(m)}$ is down-projected by $W_{\text{in}}\in\mathbb{R}^{\frac{d}{l}\times u}$, where $u\ll d$, and then quantized using an individual codebook $\mathcal{C}_{m}$.
+The selected code-vector is then up-projected by $W_{\text{out}}\in\mathbb{R}^{u\times\frac{d}{l}}$ to form ${\bm{z}}_{q_{i}}^{(m)}$:
+$\bm{z}_{e_{i}}\equiv\{\bm{z}_{e_{i}}^{(m)}\mid\bm{z}_{e_{i}}^{(m)}\in\mathbb{R}^{\frac{C_{i}H_{i}}{l}},m=1,...,l\},$ (8)
 ${\bm{z}}_{q_{i}}^{(m)}=W_{\text{out}}^{\top}\operatorname*{arg\,min}_{\bm{c}_{j}\in\mathcal{C}_{m}}||W_{\text{in}}^{\top}\bm{z}_{e_{i}}^{(m)}-\bm{c}_{j}||_{2}.$ (9)
-Here $W_{\text{in}}\in\mathbb{R}^{\frac{d}{l}\times u}$, $W_{\text{out}}\in\mathbb{R}^{u\times\frac{d}{l}}$, and $u\ll d/l$.
-The projected vector $W_{\text{in}}^{\top}\bm{z}_{e_{i}}^{(m)}$ and codebook $\mathcal{C}_{m}$ are both $l_{2}$ normalized before finding the nearest codeword.
-This equalizes the scales of encoded vectors and codewords, enhancing codebook optimization by allowing a larger subset of codewords to receive gradient updates Łańcucki et al.
+Additionally, both the projected vector $W_{\text{in}}^{\top}\bm{z}_{e_{i}}^{(m)}$ and codebook $\mathcal{C}_{m}$ are $l_{2}$ normalized before computing the distance matrix.
+This equalizes the scales of input vectors and codewords, enhancing codebook optimization by allowing a larger subset of codewords to receive gradients Łańcucki et al.
 (2020).
 Pre-training paradigm.
-Training transformers can be challenging, and joint training with VQs is even harder.
-Therefore, we propose a pre-training paradigm that includes a warm-start to enhance the learning process.
-Initially, all VQ layers are deactivated, so no quantization occurs.
-Only the encoder and decoder are updated within the CS-RVQ framework, allowing latent features to bypass the quantizers and flow directly into the decoder layers.
-We term this the "pre-training" stage.
-Once the encoder and decoder have converged by minimizing reconstruction objectives, we proceed to train the entire VQ codec as usual.
-This approach helps mitigate the distribution shift of encoder features by optimizing the encoder in advance.
-It can stabilize codebook training and improve bitrate efficiency.
-Additionally, pre-training encoders and decoders is simpler, as it avoids the quantization errors associated with VQs.
-The detailed algorithm is illustrated in Appendix A.
+Training transformers can be challenging, and jointly training them with VQ layers is even more difficult.
+To address this, we propose a pre-training paradigm that includes a warm-start to facilitate the learning process.
+Initially, all VQ layers are deactivated, meaning no quantization occurs.
+During this "pre-training" stage, only the encoder and decoder are updated within the CS-RVQ framework, allowing latent features to bypass the quantizers and flow directly into the decoder layers.
+Once the encoder and decoder have converged by minimizing reconstruction objectives, we resume training the entire VQ codec as usual.
+This approach helps mitigate the distribution shift of encoder representations by pre-optimizing the encoder.
+It helps stabilize codebook training and improve bitrate efficiency.
+Moreover, pre-training an auto-encoder is simpler, as it avoids the quantization errors associated with VQs.
+The detailed algorithm is provided in Appendix A.
 3.6 Training Objectives
-To train our codec, we employ a combination of reconstruction loss $\mathcal{L}_{recon}$ and vector quantization loss $\mathcal{L}_{vq}$.
-The reconstruction loss, $\mathcal{L}_{recon}$, consists of two components: an $l_{2}$ distance between the complex spectrum $\mathcal{X}$ and its reconstruction $\hat{\mathcal{X}}$, weighted by $\lambda_{1}$, and a multi-scale mel-spectrogram loss (Kumar et al., 2023), weighted by $\lambda_{2}$.
-These are denoted as $\mathcal{L}_{stft}$ and $\mathcal{L}_{mel}$, respectively:
+To train our codec, we use a combination of reconstruction loss $\mathcal{L}_{recon}$ and vector quantization loss $\mathcal{L}_{vq}$.
+The reconstruction loss, $\mathcal{L}_{recon}$, consists of two components: an $l_{2}$ distance between the complex spectrum $\mathcal{X}$ and its reconstruction $\hat{\mathcal{X}}$, which forces the model to reconstruct the real and imaginary parts, weighted by $\lambda_{1}$, and a multi-scale mel-spectrogram loss (Kumar et al., 2023), weighted by $\lambda_{2}$.
+These are denoted as $\mathcal{L}_{stft}$ and $\mathcal{L}_{mel}$:
 $\mathcal{L}_{recon}=\lambda_{1}\mathcal{L}_{mel}+\lambda_{2}\mathcal{L}_{stft}.$ (10)
-The vector quantization loss, $\mathcal{L}_{vq}$, comprises standard codebook and commitment losses as described in Eq.
-2.
-It is averaged within the $l$ product VQs at each bitstream and summed across $s$ transmitted bitstreams.
-The final objective for end-to-end training is the summation of $\mathcal{L}_{recon}$ and $\mathcal{L}_{vq}$.
-Figure 2: Performance evaluation results of different baseline codecs: dashed lines represent DAC baselines and solid lines represent our ESC models, with x-axis being transmission bits per second and y-axis being PESQ ()($\uparrow$)( ↑ ), Mel-Distance ()($\downarrow$)( ↓ ) and SI-SDR()($\uparrow$)( ↑ ).
+$\mathcal{L}_{vq}$ comprises the standard codebook and commitment losses as described in Equation 2.
+It is averaged across the $l$ product vector quantizers and summed over all $s$ bitstreams.
+The final objective for joint optimization is the summation of $\mathcal{L}_{recon}$ and $\mathcal{L}_{vq}$.
+To deactivate the VQ layers during the pre-training stage, $\mathcal{L}_{vq}$ is set to zero.
+Figure 2: Reconstruction quality evaluation of different baseline codecs: dashed lines represent DAC baselines and solid lines represent our ESC models, with x-axis being transmission bits per second and y-axis being PESQ ()($\uparrow$)( ↑ ), Mel-Distance ()($\downarrow$)( ↓ ) and SI-SDR()($\uparrow$)( ↑ ).
 The metrics are averaged over our composed 1158 10-second speech clips.
@@
 Datasets.
-We extract 150 hours 16kHz multilingual clean speech from the DNS Challenge 2021 dataset Reddy et al.
+We extract 150 hours of 16kHz multilingual clean speech from the DNS Challenge dataset Reddy et al.
 (2021).
-Training samples are clipped into 3-second segments, and validation samples are clipped into 10-second segments.
-For evaluation, we compile 1158 multilingual 10-second speech clips from the LibriSpeech Panayotov et al.
+Training samples are clipped into 3-second segments, and validation samples into 10-second segments.
+For evaluation, we compile 1158 multilingual 10-second speech clips with non-overlapping speakers from the LibriSpeech Panayotov et al.
 (2015), Multilingual LibriSpeech Pratap et al.
@@
 (2020) datasets.
-The speakers are non-overlapping in each split.
-Baselines. We compare our codec against the current state-of-the-art time-domain codec DAC, by reproducing three versions [^2] Reproduction settings are detailed in Appendix B.1. on our dataset:
-1) DAC-Base (Adversarial): Descript’s original released codec, operating on 16kHz audio signals. It has 74M parameter count in total.
+Baselines. We compare our ESC against the current state-of-the-art time-domain codec DAC, by reproducing three versions [^2] Reproduction settings are detailed in Appendix B.1. on our dataset:
+1) DAC-Base (adversarial): Descript’s original released codec, operating on 16kHz audio signals. It has 74M parameter count in total. Its associated discriminator has 42M additional parameter count.
 2) DAC-Tiny (adversarial): A smaller version of DAC-Base, with reduced encoder and decoder dimensions, for a fair comparison with ESC.
-3) DAC-Tiny (Non-Adversarial): A smaller and non-adversarial version of DAC to assess the impact of discriminator on improving audio fidelity.
+3) DAC-Tiny (non-adversarial): A smaller and non-adversarial version of DAC to assess the impact of discriminators on improving audio fidelity.
 Implementation details. Similar to DAC baselines, we provide different versions of ESC [^3] Complete configurations are detailed in Appendix B.2.:
-1) ESC-Base (Non-Adversarial): This base version consists of 6 encoder/decoder layers, with bitrates ranging from 1.5 to 9.0 kbps, and contains 8.39M parameters when operating at 9.0 kbps.
-2) ESC-Base (Adversarial): Trained adversarially using the same multi-scale waveform and spectrogram discriminator as DAC.
-3) ESC-Large (Non-Adversarial): A scaled-up version of ESC-Base with increased Swin Transformer layer depth, resulting in 15.58M parameters when operating at 9.0 kbps.
-The ESC variants are trained using the AdamW optimizer with a learning rate of 1e-4 and betas set to (0.9, 0.999).
+1) ESC-Base (non-adversarial): A base version codec consisting of 6 encoder/decoder layers, with bitrates ranging from 1.5 to 9.0 kbps. It contains 8.39M parameters when operating at 9.0 kbps.
+2) ESC-Base (adversarial): An adversarial version using the same multi-scale multi-band waveform and spectrogram discriminator in DAC.
+3) ESC-Large (non-adversarial): A scaled-up version with increased Swin Transformer layer depth, having 15.58M parameters at 9.0 kbps.
+Our ESC variants are trained using the AdamW optimizer Loshchilov (2017) with a learning rate of 1e-4 and a weight decay of 1e-2.
 Training runs up to 0.4 million iterations without learning rate schedulers.
@@
 (2015).
-The quantization dropout rate is set to 0.75.
+The quantization dropout rate $p$ is set to 0.75.
 Loss weighting hyperparameters are set as $\lambda_{1}=0.25$, $\lambda_{2}=1.0$, and the commitment loss weighting $\beta=0.25$.
-In ESC-Base (Adversarial), the $\mathcal{L}_{stft}$ component is replaced with the $l_{1}$ feature matching loss, following DAC’s approach.
-The adjusted loss weights are $\lambda_{1}=15.0$, $\lambda_{2}=2.0$, and $\beta=0.25$.
+For ESC-Base (adversarial), the $\mathcal{L}_{stft}$ component is eliminated.
+We use the HingeGAN Lim and Ye (2017) adversarial loss formulation and the $l_{1}$ feature matching loss Kumar et al.
+(2019), following the approach of DAC.
 Automatic evaluation metrics.
@@
 These include the PESQ score Union (2007) from the speech enhancement domain, following Jiang et al.
-(2022a), the $L_{1}$ distance between log mel-spectrograms of reference and decoded waveforms (mel distance) Kumar et al.
-(2023), and the scale-invariant source-to-distortion ratio (SI-SDR) Le Roux et al.
+(2022a); the $l_{1}$ distance between log mel-spectrograms of reference and decoded waveforms (Mel-Distance) Kumar et al.
+(2023); and the scale-invariant source-to-distortion ratio (SI-SDR) Le Roux et al.
 (2019).
-To measure codec inference latency, we use the real-time factor (RTF), defined by Défossez et al.
-(2023) as the ratio of speech audio duration to model processing time.
-4.2 Comparison with DAC
-We present a comprehensive comparison focusing on compression performance and inference efficiency, as demonstrated in Figure 2 and Table 1.
-Performance evaluation.
-First, note ESC-Base and DAC-Tiny are similar in model size, each with approximately 8 million trainable parameters.
-Our results show that ESC-Base consistently outperforms DAC-Tiny across all bitrates, even without an adversarial discriminator.
-In contrast, DAC-Tiny’s reconstruction quality drops significantly without a discriminator, particularly in SI-SDR statistics.
-This indicates a heavy reliance of DAC models on GANs for maintaining high reconstruction quality.
-Notably, ESC-Base is compatible with the same convolution-based GAN discriminator used in DAC, as evidenced by its improved performance across all metric curves in its adversarial variant.
-Besides, ESC-Large demonstrates that increasing ESC’s model size can further enhance performance, with its PESQ curve matching that of DAC-Base, the top-performing and largest model.
-While DAC-Base achieves higher SI-SDR values, ESC-Large has a smaller mel-distance.
-Thus, we conclude that the two codecs achieve comparable performance, even though ESC-Large is trained without GANs.
+To measure codec inference latency, we use the real-time factor (RTF), defined as the ratio of speech audio duration to model processing time Défossez et al.
+(2023).
 Real Time Factor $\uparrow$
@@
 Table 1: Complexity evaluation results of different baseline codecs: RTFs are measured from 100 10-second speech clips on an Intel Xeon Platinum 8352V CPU.
+4.2 Comparison with DAC
+We provide a thorough comparison focusing on compression rate, reconstruction quality, and inference efficiency, as shown in Table 1 and Figure 2.
+Performance evaluation.
+First, it is important to note that ESC-Base and DAC-Tiny are similar in model size, each with approximately 8 million trainable parameters.
+Our results show that ESC-Base consistently outperforms DAC-Tiny across all bitrates, even without an adversarial discriminator.
+In contrast, DAC-Tiny’s reconstruction quality significantly drops without a discriminator in training, particularly in SI-SDR statistics.
+This indicates a heavy reliance of DAC models on GANs for maintaining high reconstruction quality.
+Notably, ESC-Base is compatible with the same convolution-based GAN discriminator used in DAC, as evidenced by its improved performance across all metric curves in its adversarial variant.
+Additionally, ESC-Large demonstrates that increasing ESC’s model size can further enhance performance, with its PESQ curve matching that of DAC-Base, the top-performing and largest model.
+While DAC-Base achieves higher SI-SDR values, ESC-Large records a smaller Mel-Distance.
+Thus, we conclude that the two codecs achieve comparable performance, even though ESC-Large is trained without an adversarial discriminator.
 Complexity evaluation.
 Despite their exceptional performance, Descript’s top-performing codecs face significant computational challenges.
-This is evident from Table 1, where the decoding real-time factor (RTF) for DAC-Base is approximately 3.0, making it impractical for real-time applications.
-In contrast, our transformer-based ESC achieves much higher RTFs, indicating superior computational efficiency.
-Although ESC-Base is not as fast as DAC-Tiny due to the overhead of attention computation, it offers substantially better speech reconstruction capabilities, striking a favorable balance between compression performance and latency.
-Moreover, following the CS-RVQ scheme, ESC possesses enhanced encoding speeds at lower bitrates—a capability not found in DAC models.
-These results suggest that our transformer-based codec, equipped with CS-RVQ, is a more parameter-efficient foundational codec compared to time-domain convolutional codecs.
-We can conclude that ESC is a more lightweight and effective neural codec, because ESC-Large achieves comparable performance to DAC-Base without requiring a powerful GAN discriminator, with approximately $\times$4.8 smaller model size, $\times$1.4 times encoding speed, and $\times$6.4 faster decoding speed.
-4.3 Ablation Study
-To investigate the effectiveness of the proposed components in ESC, we conducted thorough ablation experiments [^4] Implementation setups are detailed in Appendix B.3.
-by training frequency-domain codecs with different architectures.
-For fair comparisons, all other ablation models listed in Table 2 have similar model sizes to ESC-Base.
-Swin Transformers and CNNs.
-To demonstrate that transformers are superior backbones for encoder/decoder in speech coding, we focus on two pairs of experiments: CNN/SwinT + RVQ and CNN/SwinT + CS-RVQ.
-In these experiments, the channel dimensions of the CNN blocks are set to match the hidden dimensions of the Swin Transformer Blocks (STBs).
-The comparison, as shown in the results, reveals that transformer-based codecs consistently outperform CNN-based codecs across all performance metrics and bitrates, regardless of the quantization scheme used.
-CS-RVQ and RVQ.
-Table 2 also highlights that CS-RVQ is a superior quantization scheme compared to RVQ, regardless of whether the backbone is CNN or STB.
-RVQ-based codecs would hit performance bottlenecks, as adding more VQs does not improve audio quality (e.g., from 6.0kbps to 9.0kbps).
-However, codecs using the CS-RVQ scheme do not face such bottlenecks at higher bitrates and consistently outperform their RVQ counterparts.
-CS-RVQ is thus a more efficient scheme that leverages multi-scale features effectively.
-Effect of pre-training paradigm.
-To evaluate the efficacy of a pre-training stage, we conducted an experiment of ESC-Base w/o pre-training.
-We monitored the VQ utilization rate, calculated as the sum of entropy (in bits) divided by the maximum number of bits from all transmitted bitstreams.
-This metric reflects bitrate efficiency and the fraction of seldom-used codewords.
-The results indicate that all experiments with pre-training achieve a near 1.0 utilization rate.
-However, ESC-Base w/o pre-training displays a lower utilization rate at 9.0kbps, and its reconstruction performance is also worse than ESC-Base.
-These findings suggest that including a pre-training stage indeed helps avoid bitrate wastage and improve audio reconstruction quality.
+This is evident from Table 1, where the decoding real-time factor (RTF) for DAC-Base is approximately 3.0, making it rather impractical for real-time applications.
+In contrast, our transformer-based ESC achieves much higher decoding RTFs (approximately 34), indicating superior computational efficiency.
+Although ESC-Base is not as fast as DAC-Tiny due to the overhead of attention computation, it offers substantially better speech reconstruction capabilities, striking a favorable balance between compression performance and computation latency.
+Future work could incorporate transformer speedup techniques such as FlashAttention Dao et al.
+(2022) to further enhance ESC’s latency further.
+Moreover, following the CS-RVQ scheme, ESC possesses faster encoding speeds at lower bitrates—a capability not evidently found in DAC models.
+These results suggest that our transformer-based codec, equipped with CS-RVQ, is a more parameter-efficient foundation model compared to time-domain convolutional counterparts.
+ESC is shown to be a more lightweight and effective neural speech codec, as ESC-Large achieves comparable performance to DAC-Base without the need for a powerful discriminator.
+Specifically, it boasts approximately $\times$4.8 smaller model size, $\times$1.4 faster encoding speed, and $\times$6.4 faster decoding speed.
 Method Bitrate PESQ $\uparrow$ Mel dist. $\downarrow$ SI-SDR $\uparrow$ VQ util. $\uparrow$
@@
 Table 2: Performance evaluation of different ablation models: results are obtained from the 1157 10-second speech clips in our test dataset.
+4.3 Ablation Study
+To investigate the effectiveness of the proposed components in ESC, we conducted thorough ablation experiments [^4] Implementation setups are detailed in Appendix B.3.
+by training frequency-domain codecs operating on complex STFT spectra with different architectures.
+For fair comparisons, all other ablation models listed in Table 2 have similar model sizes to ESC-Base.
+Swin Transformers and CNNs.
+To demonstrate that transformers are superior auto-encoder backbones in neural speech coding, we focus on two pairs of experiments: CNN/SwinT + RVQ and CNN/SwinT + CS-RVQ.
+In these experiments, the channel dimensions of the CNN blocks are set to match the hidden dimensions of the Swin Transformer Blocks (STBs).
+The comparison, as shown in Table 2, reveals that transformer-based codecs consistently outperform CNN-based codecs across all performance metrics and bitrates, regardless of the quantization scheme used.
+CS-RVQ and RVQ.
+Table 2 highlights that CS-RVQ is a superior quantization scheme compared to RVQ, regardless of whether the backbone is CNN or STB.
+RVQ-based codecs hit performance bottlenecks, as adding more VQs does not improve audio quality (e.g., from 6.0 kbps to 9.0 kbps).
+However, codecs using the CS-RVQ scheme do not face such bottlenecks at higher bitrates and consistently outperform their RVQ counterparts.
+CS-RVQ is therefore a superior vector quantization framework that leverages multi-scale features effectively.
+Effect of pre-training paradigm.
+To evaluate the efficacy of the pre-training stage, we conducted an experiment of ESC-Base w/o pre-training.
+We monitored the VQ utilization rate, calculated as the sum of entropy (in bits) divided by the maximum number of bits from all transmitted bitstreams.
+This metric reflects bitrate efficiency and the fraction of seldom-used codewords.
+The results indicate that models with pre-training achieve a near 1.0 utilization rate.
+However, ESC-Base w/o pre-training displays a lower utilization rate at 9.0 kbps, and its reconstruction performance is also inferior to that of the fully pre-trained ESC-Base.
+These findings suggest that the pre-training paradigm indeed helps avoid bitrate wastage and improve audio reconstruction quality.
 5 Conclusions
-In this paper, we introduced ESC, the first fully transformer-based neural speech codec designed for multilingual speech coding.
-ESC surpasses the existing state-of-the-art time-domain codecs in computational efficiency and achieves comparable compression performance even without a powerful adversarial discriminator.
-Our extensive evaluations demonstrate that the cross-scale residual vector quantization scheme and the Swin Transformer backbones are more efficient for neural speech coding than the traditional convolutional blocks and residual vector quantization.
-Overall, our study suggests a promising direction for speech compression.
-Future research can focus on multi-scale quantization techniques and explore additional transformer variants suitable for speech signal modeling.
+In this paper, we introduce ESC, the first fully transformer-based neural speech foundation model designed for multilingual speech coding.
+ESC surpasses existing state-of-the-art time-domain VQ-based codecs in terms of complexity and achieves comparable compression performance without the need for a powerful adversarial discriminator.
+Our extensive evaluations demonstrate that the cross-scale residual vector quantization scheme and the Swin Transformer backbones are better suited for neural speech coding than the convolutional blocks and residual vector quantization utilized in mainstream codecs.
+Overall, our study suggests a promising direction for speech foundation models.
+Future research could focus on expanding multi-scale vector quantization techniques and investigating additional transformer variants optimized for speech signal modeling.
 6 Limitations
-First, recent audio codecs are increasingly used in downstream generation tasks, where the codec acts as a foundational model to create discrete acoustic representations Borsos et al.
+First, recent neural audio codecs are increasingly utilized in downstream generation tasks, where the codec acts as a foundation model to create discrete acoustic representations Borsos et al.
 (2023); Kreuk et al.
@@
 (2024).
-These compressed representations are treated as audio tokens and can be used for language modeling, making them more suitable for generation tasks.
-Our work does not address this significant aspect.
-Therefore, a promising future direction is to re-evaluate the ESC models in similar downstream applications.
-We believe that the cross-scale representations learned from transformer backbones would offer advantages over the fixed-scale features of convolutional codecs when applied to speech synthesis tasks.
-Second, different automatic metrics for audio evaluation often produce inconsistent results, which is also evidenced in our experiments.
-To strengthen our conclusions, it is necessary to conduct subjective evaluations involving human evaluators, such as MUSHRA listening tests Series (2014).
-Third, given the scalability of transformers Kaplan et al.
-(2020), further increasing the ESC model size and training it on more extensive audio datasets is another direction to enhance compression performance.
-Lastly, as discussed in previous sections, the proposed cross-scale residual vector quantization scheme requires partial decoder layers during compression, which introduces additional latency as the bitrate increases.
-Future research could focus on refining this scheme and developing alternative multi-scale quantization structures that are more computationally efficient.
+These compressed representations, treated as acoustic tokens, are suitable for auto-regressive language modeling in generative tasks.
+However, our work does not explore this important aspect.
+A promising future direction would be to evaluate ESC in downstream applications such as speech synthesis and speech recognition.
+We anticipate that the cross-scale code representations learned from transformer backbones could offer advantages over the fixed-scale features of mainstream convolutional codecs in these tasks.
+Second, different automatic metrics for audio evaluation can produce inconsistent results, which is evidenced in our results.
+To further strengthen our conclusions, it is necessary to conduct subjective evaluations involving human evaluators, such as MUSHRA listening tests Series (2014).
+Despite this limitation, we provide a collection of demo speech samples publicly available in our codebase, which we hope will help demonstrate ESC’s performance and compensate for the absence of subjective metrics.
+Besides, the primary focus of this work is to demonstrate the superiority of transformer and cross-scale frameworks over other mainstream methods, rather than to develop a production-ready codec like DAC or EnCodec.
+Nonetheless, given the scalability of transformers Kaplan et al.
+(2020), increasing the ESC model size and training it on larger and more diverse audio datasets also represent a promising direction for enhancing its practical applicability.
+Finally, as discussed in Section 3.4, the cross-scale residual vector quantization (CS-RVQ) scheme requires the partial use of decoder layers during the encoding process, introducing additional latency as the bitrate increases.
+Similar to residual vector quantization, CS-RVQ requires careful sampling of the transmitted bitstream during training to achieve scalable bitrates within a single model.
+This sampling strategy can lead to performance trade-offs across different bitrates and may cause instability during training.
+Therefore, future research in speech foundational models could explore leveraging alternative recurrent structures Toderici et al.
+(2017); Johnston et al.
+(2018); Diao et al.
+(2020) to improve coding scalability and address these challenges.
 Appendix A Pre-training Paradigm
-The details of the proposed pre-training paradigm for efficiently learning VQ-based codecs are depicted in Algorithm 3.
-The training includes a warm start by-passing all the quantization modules and updating encoder and decoder layers only, using the reconstruction losses.
-In implementation, we set the quantization output $\bm{z}_{q}$ to be equal to the encoded representation $\bm{z}_{e}$.
-We also zero out the VQ losses to prevent codebook updates.
+The proposed pre-training paradigm for optimizing vector quantization layers is detailed in Algorithm 3.
+During the pre-training phase, all vector quantization layers are bypassed, effectively reducing the codec to a standard autoencoder trained solely on reconstruction losses (Lines 2-4).
+Once the encoder and decoder reach a certain level of convergence, the VQ layers are reactivated, and joint optimization resumes.
+In the pre-training phase, we set the $\operatorname*{arg\,min}$ nearest-neighbor selection as an identity function, making $\bm{z}_{q}$ equal to the input vector $\bm{z}_{e}$.
 Algorithm 3 Pre-training Paradigm
 1:repeat
-2:     $\hat{\mathcal{X}}=G_{\psi}(F_{\phi}(\text{STFT}(\bm{x})))$
-3:     $\mathcal{L}=\mathcal{L}_{recon}(\bm{x},\text{I-STFT}(\hat{\mathcal{X}}))$
+2:     $\hat{\mathcal{X}}=G_{\psi}(F_{\phi}(\mathcal{X}))$
+3:     $\mathcal{L}=\mathcal{L}_{recon}(\mathcal{X},\hat{\mathcal{X}})$
 4:     take gradient descent step on $\nabla_{\phi}\mathcal{L},\nabla_{\psi}\mathcal{L}$
@@
 B.1 DAC Reproduction Setups
-Our customized reproduction of DAC models largely follows their released development codes.
-The original DAC model, designed for 16kHz audio signals, uses 12 VQs, supporting bitrates from 0.5kbps to 6.0kbps.
-To compare with ESC at the same bitrate levels, we extended the number of codebooks in its RVQ to 18, creating DAC-Base.
+Our customized reproduction of DAC models closely follows the official development scripts.
+The original DAC model, designed for 16kHz audio signals, employs 12 VQ layers in its residual VQ module, supporting bitrates ranging from 0.5 kbps to 6.0 kbps.
+To ensure a fair comparison with ESC at similar bitrate levels, we extended the number of VQ layers in the RVQ module to 18, resulting in DAC-Base.
 For DAC-Tiny, we reduced the encoder dimension from 64 to 32 and the decoder dimension from 1536 to 288, while keeping other parameters unchanged.
-All these DAC baselines were trained for 0.4 million iterations with a batch size of 16 on our dataset.
-All other configurations can be referred to from the official release [^5] DAC’s official configuration is available here..
+All DAC baselines were trained for 0.4 million iterations with a batch size of 16 on our multilingual speech dataset.
+Additional configuration details can be found in the official release [^5] The official configuration for 16kHz DAC model is available at https://github.com/descriptinc/descript-audio-codec/blob/main/conf/final/16khz.yml.
 B.2 ESC Architecture Configurations
-The configurations for ESC-Base are detailed in Table 3.
-The layer depth indicates the number of STBs cascaded at each encoder/decoder layer.
+Overall, all three ESC variants are trained in distributed setups for 0.4 million iterations across 4 NVIDIA RTX 4090 GPUs with a total batch size of 36.
+These experiments took approximately 100 GPU hours.
+B.2.1 Model Parameters
+The parameter configurations for ESC-Base are provided in Table 3.
+For STFT transformation, we use a 20 ms window length and a 5 ms hop length, implemented with torchaudio.
+The number of FFT points is set to 382, resulting in a frequency dimension of 192.
+In the Swin Transformer, the layer depth represents the number of Swin Transformer Blocks (STBs) cascaded at each encoder and decoder layer.
 We use GELU activation functions and LayerNorm for normalization.
-During quantization, our codec processes two overlapped spectrum frames together (i.e., the flattened spectrum feature is reshaped from $\mathbb{R}^{W_{i}\times H_{i}C_{i}}\rightarrow\mathbb{R}^{W_{i}/2\times 2H_{i}C_{i}}$).
-Overall, the vector quantizer of ESC-Base at each bitstream uses $\log_{2}1024\times 3\times 150=4500$ bits for every 3-second input speech (1.5kbps).
-For ESC-Large, we increase the layer depth to 4 while keeping other settings the same.
-In the variant ESC-Base (Adversarial), the GAN discriminator intervenes in training only after the pre-training stage ends.
-The GAN discriminator follows the same architecture used in DAC trainings.
-All three variants are trained in distributed setups for 0.4 million iterations on 4 NVIDIA RTX 4090 GPUs with a batch size of 36, taking approximately 150 GPU hours.
-B.3 Details on Ablation Experiments
-All ablation models were trained for 0.25 million iterations.
-The ablation model SwinT+CS-RVQ is exactly our ESC-Base codec.
-The pre-training stages constitute for 0.025 million iterations.
-The Swin Transformer encoder and decoder configurations follow those in ESC-Base, as demonstrated in Table 3.
-CS-RVQ ablation codecs follow the same configuration in Table 3 as well.
-For CNN-based codecs, their channel dimensions are set to match those of SwinT-based codecs.
-Each CNN block comprises a residual unit and a downsampling/upsampling 2D convolutional layer with a stride of 2 along the frequency dimension.
-The residual unit consists of two 2D convolutional layers, with a BatchNorm layer and Parametric ReLU activations interleved between them.
-For codecs based on RVQs, we adapted the vanilla RVQs commonly used in time-domain codecs.
-To handle 2-dimensional spectrum latents, we perform RVQ within different sub-channels.
-Specifically, the flattened feature frame is first split into sub-group vectors, similar to product VQs.
-Each sub-group vector is recursively quantized, as in RVQs.
-We set the number of product VQs to 3 and the number of residual VQs to 6, ensuring the bitrate levels match those of ESC-Base (1.5kbps per bitstream, 6 in total).
-Other training configurations are the same as ESC-Base.
-All these ablation experiments together took about 100 GPU hours on 4 RTX 4090 GPUs.
+In the down-sampling/up-sampling module, we use a scaling factor of $v=2$ to un-shuffle/shuffle along the frequency resolution only.
+Before the vector quantization layers, ESC processes two overlapping time frames together.
+To implement this, the flattened spectrum feature $\mathcal{Z}$ is reshaped from $\mathbb{R}^{W_{i}\times H_{i}C_{i}}$ to $\mathbb{R}^{W_{i}/2\times 2H_{i}C_{i}}$.
+Each frame is then split into sub-vectors, down-projected, and $l_{2}$ normalized before computing the distance matrix.
+The VQ layer at each bitstream of ESC-Base consumes $\log_{2}1024\times 3\times 150=4500$ bits per 3-second input speech (i.e., 1.5 kbps bitrate).
+For the scaled-up ESC-Large variant, we increase the STB layer depth from 2 to 4 while keeping the other configurations unchanged.
 Modules Parameters Values
@@
 Vector Quantization Product VQ Size $l$ 3
-Codebook Dimension $u$ 8
+Codevector Dimension $u$ 8
 Codebook Size $K$ 1024
-Table 3: Parameter configurations of codec variant ESC-Base, which comprises 6 encoder/decoder layers.
+Table 3: Parameter configurations of model variant ESC-Base, which comprises 6 encoder/decoder layers.
+B.2.2 Adversarial Training Setup
+In the ESC-Base (adversarial) variant, the GAN discriminator is identical to the one used in DAC, consisting of a multi-period discriminator (MPD), multi-band discriminator (MBD), and multi-scale STFT discriminator (MSD), totaling over 42 million parameters.
+The adversarial loss formulation follows the official DAC-Base (adversarial) configuration.
+Additionally, we maintain the pre-training paradigm for 0.75 million iterations in this variant, with the discriminator intervening in training only after the pre-training stage finishes.
+B.3 Details on Ablation Experiments
+All ablation models operate on the complex STFT spectrum, as in ESC (SwinT + CS-RVQ), using the same STFT configurations specified in Table 3.
+These models were trained for 0.25 million iterations, with 0.025 million iterations allocated for pre-training.
+The Swin Transformer configurations mirror those used in ESC-Base.
+Similarly, the vector quantization setup in the CS-RVQ models follows that of ESC-Base.
+In total, the ablation experiments required approximately 80 hours on 4 RTX 4090 GPUs.
+B.3.1 Convolution Blocks
+For models with CNN backbones, the convolutional channel dimensions were set to match the hidden sizes of the STB-based models.
+Each CNN block consists of one residual unit and one downsampling/upsampling 2D convolutional layer with a stride of 2 along the frequency resolution only.
+The residual unit consists of two 2D convolutional layers, each followed by BatchNorm and Parametric ReLU activation.
+B.3.2 Residual Vector Quantization Setups
+For models using RVQs, we adapted the basic RVQ framework commonly used in time-domain codecs.
+To process frequency-domain spectrum features at the latent bottleneck, we combined RVQ with product vector quantization.
+Specifically, the flattened time frame vector is split into sub-group vectors, which are then recursively quantized, as in standard RVQs.
+We set the number of product VQs to 3 and the number of residual VQs to 6, ensuring the bitrate levels match those of ESC-Base (1.5 kbps per bitstream, 6 in total).